### v2

# Agent Journey Beyond RGB: Unveiling Hybrid Semantic-Spatial Environmental Representations for Vision-and-Language Navigation

Abstract
Navigating unseen environments based on natural language instructions remains difficult for egocentric agents in Vision-and-Language Navigation (VLN). While recent advancements have yielded promising outcomes, they primarily rely on RGB images for environmental representation, often overlooking the underlying semantic knowledge and spatial cues. Intuitively, humans inherently ground textual semantics within the spatial layout during indoor navigation. Inspired by this, we propose a versatile Semantic Understanding and Spatial Awareness (SUSA) architecture to facilitate navigation. SUSA includes a Textual Semantic Understanding (TSU) module, which narrows the modality gap between instructions and environments by generating and associating the descriptions of environmental landmarks in the agent’s immediate surroundings. Additionally, a Depth-based Spatial Perception (DSP) module incrementally constructs a depth exploration map, enabling a more nuanced comprehension of environmental layouts. Experimental results demonstrate that SUSA’s hybrid semantic-spatial representations effectively enhance navigation performance, setting new state-of-the-art performance across three VLN benchmarks (REVERIE, R2R, and SOON). The source code are publicly available at https://github.com/HCI-LMC/VLN-SUSA.
1 Introduction
With the advancement of multimodal technology [43, 32, 63, 47, 24], embodied VLN tasks [3, 40, 35, 37, 7, 44] have garnered significant attention due to their promising applications in robotics and intelligent assistance. Among them, conventional VLN tasks [3, 22, 20] focused on step-by-step navigation within unseen environments based on detailed instructions, while goal-oriented VLN tasks [65, 40, 45] demand agents to identify predefined objects based on high-level instructions. Consequently, agents must profoundly cognize and exploit the rich environmental information and ground textual instructions to brilliantly accomplish these arduous navigation tasks.
Figure 1: An overview of the proposed SUSA. Unlike previous approaches that solely rely on RGB input, we introduce view-level textual panoramas and trajectory-level depth exploration maps, supporting the agent’s explicit understanding of the environment.
A significant body of VLN research has focused on Transformer-based models [4, 18, 38, 55, 39] for interacting with instructions, environmental images, and historical trajectories. While these models have undergone substantial development and exhibit efficacy, two key challenges remain in VLN tasks. First, VLN agents struggle to precisely match the landmarks (e.g., “fireplace” in Fig. 1) in instructions with the entities in environments, due to the intrinsic modality heterogeneity between vision and language. To address this, recent work [11, 54, 28] has leveraged contrastive learning to promote alignment, while others have developed repositories of candidate object concepts [53, 29] or utilized large language models (LLMs) to identify the most relevant landmarks in the visual scene [30]. However, these methods often lack an explicit understanding of environments from the perspective of textual semantics (i.e., the linguistic descriptions of landmarks). Intuitively, when navigating indoors or searching for objects, humans naturally associate the textual semantics of their surroundings.
The second challenge pertains to the scarcity of diverse environmental sources, which often leads VLN agents to overfit to familiar environments, hindering generalization. To alleviate this limitation, various methods [23, 57, 17, 6, 50, 27] have employed data augmentation to increase the environmental quantity and diversity. More recent endeavors enhance spatial perception by constructing finer-grained graph [25, 2], semantic map [15, 19], grid [58, 33], or volumetric [34] representations. However, these methods predominantly rely on RGB images, overlooking the depth images, which provides more intuitive spatial structure, especially in stereoscopic navigation environments. Furthermore, few approaches [48] treat depth as a distinct modality capable of offering complementary spatial information, instead using it merely as an auxiliary signal to refine RGB-based representations. In conclusion, the co-occurrence semantic and spatial environment representations, beyond surface RGB images, has not been fully explored or exploited.
In this work, we propose the Semantic Understanding and Spatial Awareness (SUSA) architecture, designed to explicitly capture both semantic and spatial information during navigation. As illustrated in Fig. 1, we first construct a textual semantic panorama by extracting salient landmarks as the textual semantics for each view to explicitly match the textual semantics with language instructions, and select the most instruction-relevant view for navigation prediction, thereby bridging the semantic gap between the visual and linguistic modalities. To enhance the spatial awareness, we then progressively construct an independent depth exploration map, which helps mitigate the risk of the agent overfitting to the information-dense RGB images. To enabling the agent to absorb complementary information from different environmental perspectives, we adaptively fuse these hybrid semantic-spatial environmental representations to facilitate informed navigation predictions, and employ contrastive learning to align these hybrid representations with language instructions. We empirically substantiate that exclude semantic panoramas and depth exploration maps during pretraining manifest markedly enhanced generalization capabilities. Experimental results demonstrate that our agent achieves significantly improved navigation performance, outperforming state-of-the-art methods across three vision-and-language navigation benchmarks: R2R [3], REVERIE [40], and SOON [65].
The main contributions of our SUSA are as follows:
  • SUSA devises a textual-aware semantic understanding module to match the most related landmarks between environments and instructions, explicitly addressing the modality heterogeneity.
  • SUSA presents a depth-based spatial perception module to explore and represent environmental layouts beyond surface RGB content.
  • Extensive experiments on three VLN benchmarks have demonstrated the superiority of hybrid semantic-spatial environmental representations introduced by SUSA.
2 Related Work
Vision-and-Language Navigation (VLN). Vision-and-Language Navigation (VLN) tasks [3, 40, 65] require an agent to navigate sequentially towards a designated location or object based on natural language instructions. Given that the VLN task can be formulated as a partially observable Markov decision process, early methods [56, 52, 66, 13, 21, 36] primarily relied on reinforcement and imitation learning paradigms to improve navigation performance, incorporating various strategies and auxiliary loss [64, 56] designs. More recently, to better generalize to unknown environments and ground instructions, many researchers [4, 18, 38, 55, 39] have focused on transformer-based architectures, which effectively model interactions among instructions, environmental features, and historical navigation trajectories. While these approaches have propelled VLN forward, the co-occurrence environmental sources have not been fully utilized. This paper introduces text semantics and depth information to enrich environmental representations.
Semantic Alignment in VLN. To align the relationship between the environment and instructions, current researchers and practitioners in the VLN community can be divided into two main categories: implicit and explicit approaches. Implicit methods [11, 54, 28] commonly employ contrastive learning to associate visual contexts with textual instructions or use attention mechanisms to model the relationship. However, due to the inherent semantic gap between these modalities, implicit approaches have limited alignment capabilities. Explicit alignment techniques aim to project both semantics into a unified modality space. For example, some methods transform the environment into a text-based modality, Console [30] scores landmark co-occurrences to identify key navigation objects, while AACL [29] leverages CLIP [43] to refine textual entities, compacting them with visual features. However, these explicit methods often directly “stitch” textual and visual environment representations, which may leading to an overreliance on precise but abstract text while underutilizing rich yet redundant visual data. We thus use environmental textual semantic as an independent modality to explicitly match the most instruction-relevant landmarks.
Environment Representation in VLN. Limited training data and embodied environment scarcity in VLN tasks usually hinder the agent to generalize the unknown scenes or layouts. Previous approaches mitigate this issue through data augmentation [31, 57] or by enhancing visual representations[61, 34]. ScaleVLN [57] introduces thousands of photorealistic environments and FDA [17] incorporates low-frequency interference images to emphasize essential high-frequency details. Advanced visual representations, such as topological maps and grid-based layouts, further enable agents to retain temporal navigation paths and more precisely perceive spatial structures [1, 58]. For example, DUET [5] employs dual-scale transformers [51] to balance global and local decision-making, while VER [34] adopts volumetric models to represent 3D spatial structures. While most research relies on RGB representations, which tend to capture redundant appearance features, our method incorporates depth data, offering agents a more intuitive and spatially aware understanding of the environment.
Figure 2: The detailed architecture of our proposed SUSA model. The orange and green arrows highlight our proposed Textual-Aware Semantic Understanding (TSU) and Depth-Aware Spatial Perception (DSP) modules, respectively. The Hybrid Representation Fusion (HRF) module is designed for aligning environmental representations with instructions.
3 Methodology
3.1 Preliminary
Task Formulation. We focus on discrete indoor VLN tasks [3, 40, 65] within a Matterport3D simulator [3], modeled as an undirected graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}$ denotes nodes and $\mathcal{E}$ signifies the connectivity paths. At the outset of each episode, the agent, equipped with a GPS sensor, RGB and depth cameras, is situated at a randomly navigable node and receives a language instruction $I=\{\text{w}_{i}\}_{i=1}^{l}$ including $l$ words. We use BERT [9] to extract instruction features and then embed them as =$\mathcal{I}=\{w_{1},\dotso,w_{l}\}$. At each time step $t$, the agent can perceive the surrounding RGB panorama $R_{t}=\{\text{r}_{t,i}\}_{i=1}^{n=36}$ and depth panorama $D_{t}=\{\text{d}_{t,i}\}_{i=1}^{n=36}$ along with $n$ corresponding views. During navigation, the agent must decide its next action $a_{t}$ by selecting one navigable node from the candidate actions $\mathcal{A}_{t}=\{a_{t,i}\}_{i=0}^{k}$. The sequential decision-making process is iterated until the agent decides to stop at the current location or exceeds the maximum navigation steps. For goal-oriented tasks [40, 65], the agent is also tasked with locating a target object.
Overview of the Proposed Method. Our objective is to equip the agent with a textual semantic panorama and a depth exploration map to accord complementary environmental content. We adopt DUET [5] as our baseline, which merely utilizes RGB panoramas and topological maps as environmental representations for both local and global action predictions. The SUSA framework consists of three main modules: (a) textual-aware semantic selection is detailed in Sec. 3.2, (b) depth-aware visual environment exploration is presented in Sec. 3.3, and (c) hybrid representation fusion and action prediction is subsequently introduced in Sec. 3.4, as illustrated in Fig. 2.
3.2 Textual-Aware Semantic Understanding
In this module, we build a textual semantic panorama that enables the agent to understand the environment from a textual perspective, allowing it to select the most relevant navigable view as the optimal local navigation action.
Figure 3: Illustration of TSU module in  Fig. 2, which matches instructions and semantic features by static and dynamic strategies.
Salient Landmarks Extractor. The pretrained visual-language model readily provides the capability for agents to perceive textual semantics by generating detailed captions from RGB images. As illustrated in Fig. 2, we utilize off-the-shelf BLIP-2-FlanT5-xxL [24], a powerful visual-language model, to discretize $R_{t}$ and generate corresponding textual descriptions for each view using the simple prompt “a photo of”. Following [14, 23], this process yields a textual panorama $\mathcal{S}_{t}=\{s_{t,i}\}_{i=1}^{n=36}$, where each view $s_{t,i}$ contains informed descriptions of salient landmarks. These descriptions typically involve room types and objects (e.g.,“a living room with two chairs”), which assist in the object recognition. We further discuss the quality of the generated descriptions in supplementary materials. Additionally, we employ the CLIP (ViT-L-14-336px) text encoder [43] to extract textual semantic panorama features, embedding them as $\mathcal{X}^{\mathcal{S}}_{t}=\{x^{s}_{t,1},\dots,x^{s}_{t,n}\}$. During navigation, we leverage the frozen BLIP-2 and CLIP models to empower the agent to obtain real-time textual semantic features from the navigable nodes in its vicinity.
Textual Semantic Selector. Previous methods [53, 29] aligned environmental RGB images with instructions by exploiting textual semantics yet the inherent modality discrepancy between visual and language modalities remained unaddressed. In contrast, we treat the textual semantic as a distinct modality, directly matching salient landmarks with the instructions. As shown in Fig. 2 and Fig. 3, our textual semantic selector combines both $static$ and $dynamic$ matching strategies to select the most instruction-relevant view in the textual panorama $\mathcal{X}^{\mathcal{S}}_{t}$.
Static Matching: The most straightforward approach to semantic alignment is by calculating similarity, where higher similarity indicates stronger relevance between the textual semantics of each view and the instruction. Consequently, we calculate the cosine similarity matrix $\mathbf{M}\in\mathbb{R}^{n\times l}$, which represents the correlation between each word in the instruction $\mathcal{I}$ and each view in the textual panorama $\mathcal{X}^{\mathcal{S}}_{t}$. Nevertheless, the textual semantics generated only correspond to salient landmarks (e.g.,“plant,” or “table”) rather than actions (e.g., “go to” or “ turn left”) in the instructions. Therefore, we then apply row-wise max-pooling to purify the static instruction relevance $\Gamma_{t}^{stat}\in\mathbb{R}^{n}$ to match the most relevant landmarks between textual semantics and instructions at step $t$ :
$\small\begin{split}\mathbf{M}_{i,j}=\frac{\mathcal{X}^{\mathcal{S}}_{t,i}\cdot\mathcal{I}_{j}}{\|\mathcal{X}^{\mathcal{S}}_{t,i}\|\|\mathcal{I}_{j}\|},\ \ \ \ \Gamma_{t}^{stat}=\max_{j}\mathbf{M}_{i,j}\end{split}$ (1)
where $i=1,\dots,n$ and $j=1,\dots,l$. Although straightforward, the static strategy relies on the monotonic similarity matrix to match textual semantics $\mathcal{X}^{\mathcal{S}}_{t}$ and instructions $\mathcal{I}$, struggling with long-term navigation reasoning.
Dynamic Matching: Furthermore, a dynamic matching strategy is introduced to better capture the intricate relationships between the environment and instruction. Specifically, we utilize a standard multi-layer transformer decoder to dynamically match the relevance $\Gamma_{t}^{dyn}$ between textual semantics and instructions, as formalized by:
$\small\begin{split}\Gamma_{t}^{dyn}=\text{TCA}(\mathcal{S}_{t},\mathcal{I},\mathcal{I}),\end{split}$ (2)
where TCA denotes a textual modality cross-attention mechanism that models textual semantic-instruction relationships. TCA comprises a cross-attention layer, a feed-forward network (FFN) layer, and LayerNorm, enabling the agent to adapt to long-term navigation sequences.
To trade off computational efficiency and navigation accuracy, we flexibly combine the relevance of static $\Gamma_{t}^{sim}$ and dynamic $\Gamma_{t}^{att}$. The latent semantic panoramic feature $\tilde{\mathcal{X}_{t}^{S}}$ is expressed as follows:
$\begin{split}\tilde{\mathcal{X}_{t}^{S}}=\delta\ast\Gamma_{t}^{stat}+(1-\delta)\ast\Gamma_{t}^{dyn},\end{split}$ (3)
where the balance factor $\delta$ regulates the relative contributions of each matching strategy. By mapping the environment to the same textual modality, the textual semantic selector enables the agent to better understand the environment context from a textual semantic perspective, thus facilitating grounded language instructions.
3.3 Depth-Aware Spatial Perception
Depth images, which also contain more intuitive and easily distinguishable structural information, were not adequately exploited. To enhance the spatial perception for agent, we concurrently construct a depth-based exploration map to memorize depth-based navigation trajectory alongside the RGB-based one [5].
Depth-Based Exploration Mapping. Concretely, we obtain ground truth depth images for each view from the Matterport3D simulator. As illustrated in Fig. 2, we then employ CLIP (ViT-L/14@336px) and ResNet-50 [16] (pretrained on the Gibson [59]) as the depth and RGB encoders, respectively, to extract panoramic depth features $\mathcal{D}_{t}\in\mathbb{R}^{n\times d_{v}}$ and panoramic RGB features $\mathcal{R}_{t}\in\mathbb{R}^{n\times d_{v}}$, where $d_{v}$ represents the dimension of each view. Moreover, we employ a two-layer transformer-based panorama encoding module to perform self-attention and sequentially encode the depth and RGB panoramic images. The panorama encoding modules for both modalities share weights, since each view in the depth and RGB panoramas is accordant. Subsequently, we aggregate each visited node and its adjacent views by average pooling to construct depth and RGB exploration maps in parallel. The explored depth and RGB trajectory maps are formalized as $\mathcal{T}^{\mathcal{D}}\in\mathbb{R}^{k\times d_{n}}$ and $\mathcal{T}^{\mathcal{R}}_{t}\in\mathbb{R}^{k\times d_{n}}_{t}$ Here, $k$ denotes the number of navigable nodes at step $t$, while $d_{n}$ represents the dimension of each navigable node. These exploration maps memorize the agent navigation trajectory information for global action predictions, which serves to guide the agent traversal across various navigable points when considering potential backtracking.
Cross-Modal Interaction and Reasoning. To defer to instructions, we utilize a multi-layer cross-encoder that interacts with both the language instructions and the depth exploration map features to generate the refined depth map $\tilde{\mathcal{T}^{\mathcal{D}}}$. This encoder comprises four residual connection layers of a cross-modal transformer, each incorporating graph-aware self-attention [5] (GASA), cross-modal attention, and a feed-forward network. The GASA operation is formally expressed as $\text{GASA}(X)=\eth(X\Theta_{q}(X\Theta_{k})^{T}+\mathcal{M})X\Theta_{v}$ where $\mathcal{M}$ represents a learnable distance matrix encoding the relative distances between navigation points, $\eth$ denotes the Softmax activation function, and $\Theta_{q},\Theta_{k},\Theta_{v}$ are learnable weight matrices. In parallel, we facilitate similar cross-modal interactions between the language instructions and the RGB exploration map $\mathcal{T}^{\mathcal{R}}$, as well as the RGB panoramas $\mathcal{X}^{\mathcal{R}}$. These are encoded as $\tilde{\mathcal{T}^{\mathcal{R}}}$ and $\tilde{\mathcal{X}^{\mathcal{R}}}$, thereby enabling the model to perform both global and local predictions based on the given instructions $\mathcal{I}$. Compared to the standalone $\mathcal{T}^{\mathcal{R}}$, the incorporation of a depth exploration map $\mathcal{T}^{\mathcal{D}}$ enhances the perceptual clarity of environmental structures, mitigating the risk of overfitting to RGB-based visual noises and biases.
Figure 4: Illustration of the proposed HRF module in  Fig. 2.
3.4 Hybrid Representation Fusion and Prediction
Building on aforementioned methods, at the current state $t$, the agent synchronously perceives the environment through multiple modalities—textual semantic, depth, and RGB —at various scales (local and global). To encourage the agent to aggregate complementary context from these hybrid representations, we encode these features into a low-dimensional latent space via attention pooling and apply contrastive learning to align them with the instructions.
Fusing Hybrid Semantic-Spatial Environmental Representations. As depicted in Fig. 4, we employ attention pooling to project all environmental representations at time step $t$, along with the corresponding instruction, into an equivalent dimension $d$. Specifically, the attention pooling AP operation is defined as follows:
^$\hat{\mathcal{H}}=\text{AP}(\mathcal{H},v),\quad\mathcal{H}\in\{\mathcal{X}^{\mathcal{S}},\mathcal{T}^{\mathcal{D}},\mathcal{X}^{R},\mathcal{T}^{R},\mathcal{I}\},$ (4)
where $v\in\mathbb{R}^{d}$ denoting the learnable query vector and ^$\hat{\mathcal{H}}\in\mathbb{R}^{d}$. The AP operator is given by:
$\eta=\eth\left(\tau(\mathcal{H})\otimes v^{\top}\right),\quad\hat{\mathcal{H}}=\tau\left(\eta\odot\mathcal{H}\right),$ (5)
with $\tau$ representing the tanh activation function. Subsequently, we perform weighted fusion of the hybrid environmental representations $E^{\mathrm{hyb}}$, as follows:
$\small\footnotesize\mathcal{B}=\begin{bmatrix}\beta_{1}&\beta_{2}&\beta_{3}&\beta_{4}\end{bmatrix}^{\top}=\sigma(\mathrm{FFN}[\mathcal{X}_{0}^{\mathcal{S}};\mathcal{X}_{0}^{\mathcal{R}};\mathcal{T}_{0}^{\mathcal{D}};\mathcal{T}_{0}^{\mathcal{R}}])$ (6)
$\small\footnotesize E^{\mathrm{hyb}}=\begin{bmatrix}\hat{\mathcal{X}^{\mathcal{S}}}&\hat{\mathcal{X}^{\mathcal{R}}}&\hat{\mathcal{T}^{\mathcal{D}}}&\hat{\mathcal{T}^{\mathcal{R}}}\end{bmatrix}\cdot\mathcal{B}$ (7)
where $\sigma$ stands for the sigmoid activation function, and [;][;][ ; ] denotes concatenation.
[…]
$P_{\mathcal{X}}=\beta_{1}P_{\mathcal{X}^{\mathcal{S}}}+\beta_{2}P_{\mathcal{X}^{\mathcal{R}}},$ (8)
$P_{\mathcal{T}}=\beta_{3}P_{\mathcal{T}^{\mathcal{D}}}+\beta_{4}P_{\mathcal{T}^{\mathcal{R}}}.$ (9)
We convert local predictions into the global action space and predict action $a_{t}$ via dynamic fusing [5] $P_{\mathcal{T}}$ and $P_{\mathcal{X}}$.
3.5 Training and Inference
Pretraining. Our pre-training phase followed prior work [5, 25], utilizing auxiliary tasks such as Masked Language Modeling (MLM), Masked Region Classification (MRC), and Single-step Action Prediction (SAP). Object Grounding (OG) task to localize target objects in REVERIE and SOON. Given the sparsity of depth images and textual semantics compared to RGB images, completely pre-train the model may lead to overfitting and increased training costs. As shown in Fig. 2, only the computation flows indicated by black arrows were pre-trained. The effectiveness of this partial pre-trained strategy is demonstrated in Sec. 4.3.
Methods Validation Seen Validation Unseen Test Unseen
[…]
RCM [56] 10.70 23.33 21.82 16.23 15.36 11.98 9.29 6.97 4.89 3.89 10.60 7.84 6.67 3.67 3.14
HOP+ [42] 10.59 55.87 49.55 40.76 36.22 14.57 36.07 31.13 22.49 19.33 15.17 33.82 28.24 20.20 16.86
DUET [5] 13.86 71.75 63.94 57.41 51.14 22.11 46.98 33.73 32.15 23.03 21.30 52.51 36.06 31.88 22.06
DSRG [53] - 75.69 68.09 61.07 54.72 - 47.83 34.02 32.69 23.37 - 54.04 37.09 32.49 22.18
GridMM [58] - - - - - 23.20 51.37 36.47 34.57 24.56 19.97 53.13 36.60 34.87 23.45
AZHP [60] 13.60 70.98 62.24 56.99 50.14 22.08 49.02 36.25 32.41 24.13 21.10 52.52 36.11 32.10 22.54
FDA [17] - - - - - 19.04 47.57 35.90 32.06 24.31 17.30 49.62 36.45 30.34 22.08
CONSOLE [30] - 74.14 65.15 60.08 52.69 - 50.07 34.40 34.05 23.33 - 55.13 37.13 33.18 22.25
KERM [26] 14.25 71.89 64.04 57.55 51.22 21.85 49.02 34.83 33.97 24.14 18.38 52.26 37.46 32.69 23.15
VER [34] 16.13 75.83 66.19 61.71 56.20 23.03 55.98 39.66 33.71 23.70 24.74 56.82 38.76 33.88 23.19
SUSA (Ours) 14.60 76.95 69.07 61.77 55.86 22.59 51.75 38.86 35.02 26.56 17.86 54.39 41.54 36.11 27.31
Table 1: Comparison with the state of the art on REVERIE. Bold highlight the best performance in each column, $\uparrow$ indicates better performance with higher values.‘–’: unavailable statistics.
Fine-tuning and Inference. During fine-tuning, we employ a contrastive learning loss, $\mathcal{L}_{cl}$ to align the hybrid environmental representations $E^{hyb}$ with the corresponding instructions $\mathcal{I}$:
$\textstyle\mathcal{L}_{cl}=-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{j})/t)}{\sum_{j=1}^{B}\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{j})/t)}$ (10)
$\textstyle-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((\mathcal{I}_{i},E^{\mathrm{hyb}}_{j})/t)}{\sum_{j=1}^{B}\exp((\mathcal{I}_{i},E^{\mathrm{hyb}}_{j})/t)},$
where $B$ and $t$ represent the batch size and the temperature. Then, we train the model based on hybrid environment representations by performing single-step action prediction [4, 5] with a contrastive learning loss $\mathcal{L}_{cl}$. The hybrid single action prediction loss in behavior cloning is computed as $\mathcal{L}_{hsap}^{*}=\sum_{t=1}^{T}-\log p(a_{t}^{*}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$ and $\mathcal{L}_{hsap}^{\star}=\sum_{t=1}^{T}-\log p(a_{t}^{\star}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$, where $a_{t}^{*}$ represents the ground truth action, and $a_{t}^{\star}$ denotes the pseudo-target action sampled from current exploration map. The final training loss is then expressed as follows:
$\mathcal{L}_{SUSA}=\lambda_{1}\mathcal{L}_{hsap}^{*}+\mathcal{L}_{hsap}^{\star}+\lambda_{2}\mathcal{L}_{cl}$ (11)
where $\lambda_{1}$ and $\lambda_{2}$ denote loss weights.
During the inference phase, the agent is tasked with sequentially navigating to the predicted navigable points, following the shortest path algorithm [12], without making any abrupt “jumps” between points.
4 Experiments
4.1 Task Setup and Implementation
Datasets. We evaluate our model on three diverse VLN benchmarks based on the Matterport3D simulator environment [3]. R2R [3] focuses solely on navigation following detailed instructions. REVERIE [40] requires the agent to recognize the correct object from candidate bounding boxes upon reaching the navigation goal. SOON [65] further challenges agents to generate candidate bounding boxes using an object detector, as predefined boxes are not provided.
Metrics. We evaluate the navigation performance using standard metrics: Trajectory Length (TL), Success Rate (SR), Navigation Error (NE), and Success Rate weighted by Path Length (SPL). Additionally, we use Normalized Dynamic Time Warping (nDTW) and Success weighted by Normalized Dynamic Time Warping (sDTW) to evaluate the fidelity of successful trajectories to their corresponding instructions. For the REVERIE and SOON tasks, which involve object identification, we also evaluate object grounding metrics: Remote Grounding Success (RGS), and RGS weighted by Path Length (RGSPL).
Implementation Details. We initialized transformer layers with pretrained LXMERT [49] during pre-training. Pre-training was performed for 100k iterations with a batch size of 32. Fine-tuning was performed for 25k iterations. For fine-tuning, we used batch sizes of 4, 8, and 2 for R2R, REVERIE, and SOON, respectively. The default hyperparameters are $\lambda_{1}$ = 0.2 and $\lambda_{2}$= 0.8. The node and view dimensions are identical: $d_{v}=d_{n}=d=768$. All experiments were conducted on a single NVIDIA RTX 4090 GPU. We minimized changes to the baseline [5] settings and used no additional annotations. Detailed experimental setup is provided in the supplementary material.
4.2 Comparisons with State of the Art
We conduct a comprehensive comparison on the REVERIE [40], R2R [3], and SOON [65] datasets. For fairness, we exclude methods that involve pre-exploration [66, 46] or utilize large-scale data augmentation [6]. REVERIE. Table 1 presents a comparative analysis on the REVERIE. On the test unseen split, SUSA achieves an SPL score of 41.54%, significantly outperforming the previous state-of-the-art (SOTA) method by a large margin of 2.78%. Furthermore, we establish new SOTA results for RGS and RGSPL on both the validation and test unseen splits, demonstrating that the introduced textual semantics substantially elevate overall navigation, particularly benefiting object localization performance. R2R. The compared experiment results on R2R, summarized in  Tab. 2. Compared with the SOTA model GOAT [54] on the validation and test unseen splits, SUSA achieves 64.85% and 63.83% on SPL, respectively. These gains manifest not only an improved success rate but also enhanced navigation efficiency, since SPL considers both success rate and trajectory length.
Methods Validation Unseen Test Unseen
[…]
RCM [56] 11.46 43 - 6.09 11.97 43 38 6.12
HOP [41] 12.27 64 57 3.80 12.68 64 59 3.83
DUET [5] 13.41 71 61 3.30 14.73 69 59 3.65
DSRG [53] - 73 62 3.00 - 72 61 3.33
AZHP [60] 13.68 71 60 3.25 14.47 69 59 3.43
PanoGen [23] 13.40 74 64 3.03 14.38 71 61 3.31
GridMM [58] - 75 64 - - 73 62 -
FDA [17] 13.68 72 64 3.41 14.76 69 62 3.41
CONSOLE [30] 13.59 73 63 3.00 14.31 72 61 3.30
KERM [26] 13.54 71.95 60.91 3.22 14.60 69.73 59.25 3.61
ESCENE [62] 10.80 68 64 3.39 11.89 66 63 3.77
GOAT* [54] 13.52 73.95 63.36 2.82 14.91 72.83 62.42 3.29
SUSA (Ours) 12.18 73.05 64.85 3.06 13.27 72.56 63.83 3.23
Table 2: Comparison with state of the art on the R2R dataset. Bold and Underline highlight the best and the runner-up performance in each column. $\downarrow$ indicates better performance with lower values.* expressing our reimplementation results.
SOON. On the SOON dataset, as evidenced in  Tab. 3, SUSA outperforms the SOTA model GOAT by 2.27% and 0.29% on the validation and test unseen splits, respectively, in terms of SPL. Overall, the outstanding experimental results show that textual semantics boost navigation success via salient landmark matching, while the depth exploration map is beneficial to generalizable navigation efficiency.
Method Validation Unseen Test Unseen
[…]
GBE [65] 28.54 19.52 13.34 1.16 21.45 12.90 9.23 0.45
DUET [5] 36.20 36.28 22.58 3.75 41.83 33.44 21.42 4.17
GridMM [58] 38.92 37.46 24.81 3.91 46.20 36.27 21.25 4.15
KERM [26] 35.83 38.05 23.16 4.40 - - - -
GOAT[54] - 40.35 28.05 5.85 - 40.50 25.18 6.10
SUSA(Ours) 36.86 42.89 30.82 6.48 36.13 36.87 25.47 5.93
Table 3: Comparison with the state of the art on the SOON.
4.3 Diagnostic Experiment
1) Overall Design.
[…]
1 13.41 3.10 79.86 71.43 60.63
2 13.85 3.08 80.42 72.80 62.71
3 13.78 3.06 81.65 72.92 63.16
4 12.18 3.06 79.99 73.05 64.85
5 14.58 3.08 82.08 73.95 62.84
Table 4: Ablation study on different components of SUSA on the R2R validation unseen split.
To thoroughly evaluate the efficacy of key components, we conduct ablation experiments on the R2R validation unseen split in the Tab. 4. As shown in (#2, #3, #4), both the Textual-aware Semantic Understanding (TSU) and Depth-aware Spatial Perception (DSP) modules consistently outperform the baseline (#1) [5], underscoring the substantial performance gains achieved by incorporating textual panoramas and depth exploration maps over relying solely on RGB. The hybrid representation fusion (HRF) module improves the OSR/SR score from 73.05%/79.99% to 73.95%/82.08%, demonstrating that the proposed HRF enables more precise alignment of environment-instruction pairs through contrastive learning. Analysis of the HRF module and the computational efficiency of the entire model is presented in supplementary materials.
[…]
ViT No 51.07 46.98 33.73 32.15 23.03
CLIP No 57.17 51.55 35.75 35.19 24.40
CLIP ResNet(ImageNet) 54.42 50.18 34.63 34.39 23.66
CLIP ResNet(Gibson) 56.72 52.00 39.21 35.05 26.54
Table 5: Ablation of different visual representations. Gibson and ImageNet serve as pre-training sources for the depth encoder.
$\delta$ REVERIE R2R
[…]
$w/o$ 52.00 39.21 35.05 26.54 72.80 62.71 69.33 60.21
0 53.91 38.87 36.84 26.82 72.20 63.78 70.74 61.46
0.5 55.07 39.50 37.12 26.70 73.05 64.85 71.05 62.17
1.0 53.51 37.86 36.69 25.92 72.37 62.95 69.17 60.55
$adaptive$ 52.43 38.22 36.04 26.24 71.95 63.10 69.98 60.42
Table 6: Ablation on the relative importance of static and dynamic matching in the TSU module, balanced by the balance factor $\delta$.
2) Image Features vs. Structural Features. We analyze various visual features (depth and RGB) contributions within the DSP module (Tab. 5). For RGB images, substituting ViT [10] features with those from CLIP [43] yields marked improvements across all metrics, underscoring that CLIP achieves a better understanding of the environment than its ViT counterpart. As for depth features, we leveraged depth features extracted by ResNet-50 [16] pre-trained on the ImageNet [8] or Gibson [59] datasets to investigated the effect of depth exploration map. Compared to Matterport3D, the Gibson simulator offers more detailed and realistic physical environments along with higher-resolution visual information. The results demonstrate that simply incorporating depth information does not necessarily lead to improved performance. In comparison, the ResNet-50 pre-trained on the Gibson dataset extracts structural features that refine spatial perception, yielding notable improvements in the SPL metric without compromising SR or RGS. These improvements indicate that better spatial perception enhances navigation efficiency.
3) Static Matching vs. Dynamic Matching. To assess the impact of static matching ($\delta=0$) and dynamic matching ($\delta=1$) strategies in the TSU module, we perform ablation studies based on balance factor $\delta$ on the R2R and REVERIE.  Tab. 6 summarizes these results, where w/o indicates the exclusion of the TSU module, corresponding to the results in Tab. 4 (#2) and last row in Tab. 5. Furthermore, adaptive denotes that $\delta$ is a learnable parameter which adaptively adjusts the contribution of each matching strategy. Compared to the first row in Tab. 6, overall, incorporating the TSU module (regardless of static or dynamic matching strategy) improves SR and RGS, albeit at the cost of SPL on REVERIE. In contrast, it enhances SPL on R2R. While the TSU module can achieve superior performance under the adaptive setting, it falls short when $\delta$ = 0.5. Specifically, at $\delta$ = 0.5, the SPL metric reaches 39.50% and 64.85% on the two datasets, with the RGS, nDTW, and sDTW metrics also achieving optimal performance. Excellent results demonstrate dual merits of proposed TSU module, which boosts object identification accuracy on the REVERIE task while simultaneously improving the fidelity between navigation trajectories and instructions on the R2R.
Figure 5: a) and b) show the performance gap ($\downarrow$) between seen and unseen environments, while c) and d) present key metrics for different pretraining strategies on the R2R and REVERIE.
4) Improved Generalization via Pretraining. As analyzed before, pretraining the entire SUSA framework on multiple auxiliary tasks may lead to overfitting, which can negatively impact the generalization in unseen environments, particularly considering the modality differences between textual panorama, depth exploration maps, and RGB representations. As shown in Fig. 5, we observe that partial pretraining strategies not only consistently narrow the performance gap between seen and unseen environments but also outperform the completely/random pretraining strategies in navigation performance. For example, as depicted in Fig. 5 (a), on the REVERIE, the partial pretraining model achieves an SR gap of only 21.11% between the seen and unseen splits, reducing the gap by 5.92% compared to the complete pretraining strategy. Additionally, as shown in Fig. 5 (c), multiple metrics indicate that partial pretraining results in superior performance. Fig. 5(b) and (d) illustrate similar trends are observed on R2R, further emphasizing the impressive generalization power of the partial pretraining strategy.
4.4 Qualitative Analysis
As shown in  Fig. 6, compared to DUET, our SUSA agent is able to accurately stop near the “living room with a couch.” In contrast, the DUET agent, which relies exclusively on RGB environment representations, is prone to making elusive actions, hindering its ability to accurately follow instructions and reach the target location. This manifest that, thanks to the salient landmarks provided by the SUSA architecture, the agent is better able to navigate to the target location by grounding textual instructions.
Figure 6: Predicted trajectories of SUSA and DEUT on the R2R validation unseen split.
5 Conclusion
We present the SUSA architecture, which offers complementary environmental representations beyond RGB images. We propose the TSU module, enabling the agent to identify the most instruction-relevant views through both static and dynamic matching. The DSP module enhances spatial awareness by integrating depth exploration maps. We further leverage contrastive learning to precisely align hybrid environmental representations with instructions. Experimental results on the R2R, REVERIE, and SOON benchmarks demonstrate the effectiveness and generalization of our approach. Future work will enrich beneficial environmental representations to advance specific VLN tasks.
Supplementary Material
6 Detailed Experimental Setups
Metrics. We categorize the metrics into three types: Navigation, Objects Grounding, and Instruction Following. Detailed calculations of these metrics can be found in  Tab. 7. For Navigation metrics, Trajectory Length (TL): Average path length in meters. It is worth noting that since the trajectory needs to match the ground truth trajectory described by the instruction, strictly speaking, the trajectory length is not necessarily the smaller, the better. Success Rate (SR): Proportion of paths where the agent reaches within 3 meters of the target location. Oracle Success Rate (OSR): The proportion of trajectories where at least one node is within 3 meters of a corresponding node in the reference trajectory. Navigation Error (NE): Average final distance (meters) between the agent and the target. Success Rate weighted by Path Length (SPL): Prioritizes success for shorter paths, which normalizes the success rate by trajectory length. For Instruction Following metrics, we apply Normalized Dynamic Time Warping (nDTW) and Success Rate Weighted by Dynamic Time Warping (sDTW). The former measures how well the VLN agent could follow the instruction, while the latter considers nDTW for success cases. For Objects Grounding metrics: Remote Grounding Success (RGS): Proportion of instructions where the agent correctly identifies the target object. RGS weighted by Path Length (RGSPL): Prioritizes correct object identification for shorter paths.
Metric $\uparrow$ $\downarrow$ Definition Dataset
         Navigation:
TL $\sum_{1\leq i<|P|}d(p_{i},p_{i+1})$ R2R REVERIE SOON
NE $\downarrow$ $d(p_{|P|},g_{|G|})$
ONE $\downarrow$ $\min_{p\in P}d(p,g_{|G|})$
SR $\uparrow$ $\mathbb{I}[NE(P,G)\leq d_{th}]$
OSR $\uparrow$ $\mathbb{I}[\mathrm{ONE}(P,G)\leq d_{lh}]$
SPL $\uparrow$ $\mathrm{SR}(P,G)\cdot\frac{d(p_{1},g_{|G|})}{\max\{\mathrm{TL}(P),d(p_{1},g_{|G|})\}}$
         Instruction Following:
nDTW $\uparrow$ $\exp\left(-\frac{\min_{W\in\mathcal{W}}\sum_{(i_{k},j_{k})\in W}d(p_{i_{k}},q_{j_{k}})}{|P|\cdot d_{th}}\right)$ R2R
sDTW $\uparrow$ $SR(P,G)*nDTW(P,G)$
         Objects Grounding:
RGS $\uparrow$ $\mathbb{I}\Big{(}\max_{j}IoU(B_{i}^{p},B_{j}^{g})\geq 0.5\Big{)}$ REVERIE SOON
RGSPL $\uparrow$ $\mathrm{RGS}(P,G)\cdot\frac{d(p_{1},g_{|G|})}{\max\{\mathrm{TL}(P),d(p_{1},g_{|G|})\}}$
Table 7: Common VLN metrics, categorized into three types: Navigation, Instruction Following, and Object Grounding. The symbols $\uparrow$ and $\downarrow$ signify if a higher or lower metric value is preferable, respectively. P represents the paths taken by the agent, while G stands for the ground truth trajectories. The function $\mathbb{I}(\cdot)$ serves as an indicator, with $d(n,P)$ measuring the distance from node $n$ to path $P$ , and $d_{th}$ indicating the distance threshold. $W=w_{1..|W|}$ is a warping with $w_{k}=(i_{k},j_{k})\in[1:|P|],[1:|G|]$, respecting the step-size.$B^{p}_{i}$ means the i-th object’s bounding box in agent path (or trajectory).
Pre-training Tasks. Following prior works [5, 4, 25], during pre-training, we employ a suite of multimodal pre-training tasks, including MLM (Masked Language Modeling), MRC (Masked Region Classification), SAP (Single Action Prediction), and OG (Object Grounding). OG is specifically designed for object grounding using the REVERIE and SOON datasets. MLM and MRC aim to establish a robust vision-language alignment. SAP enables the agent to learn preliminary navigation knowledge by predicting the next action during pre-training. However, as navigation is inherently a sequential decision-making process, SAP can only learn static, single-step decisions.
Dataset Train Validation Seen Validation Unseen Test Unseen VLN task
instr house instr house instr house instr house
R2R [3] 14039 61 1021 56 2349 11 4173 18 Fine-grained
REVERIE [40] 10466 60 1423 46 3521 10 6292 16 Goal-oriented
SOON [65] 2779 34 113 2 339 5 615 14 Goal-oriented
Table 8: Comparison of the three VLN datasets utilized in this study.
Dataset Details. We evaluate the navigation capabilities of the agent on three distinct VLN datasets: the fine-grained R2R [3], as well as the goal-oriented REVERIE [40] and SOON [65] datasets, as shown in Tab. 8. Each dataset is divided into train, validation seen, validation unseen, and test unseen splits. REVERIE contains approximately 20k high-level instructions describing target locations and objects, with an average instruction length of 21 words. Given predefined object bounding boxes within each panorama, the agent identifies the correct bounding box upon reaching the navigation goal. Unlike the R2R dataset, in the REVERIE task, the agent must identify the correct bounding box from the candidates once the navigation target is reached. R2R includes 22k step-by-step instructions, with an average instruction length of 32 words. The R2R dataset focuses solely on the navigation task, providing detailed path planning for the agent in the instructions. SOON is another goal-oriented dataset, containing about 5k instructions, each averaging 47 words in length.
We evaluated the experimental results of the unseen test split on the public online leaderboards***https://eval.ai/web/challenges/challenge-page/606/participate, https://eval.ai/web/challenges/challenge-page/97/participate. We assure the readers that we will make our code and model checkpoints publicly available at https://github.com/HCI-LMC/VLN-SUSA.
7 Additional Experiments
7.1 Comparison with Previous Approaches
In this subsection, we discuss the differences between the proposed SUSA architecture and previous approaches incorporating depth or textual environmental information. For depth-based methods, such as BEVBert [1] and GridMM [58], depth information is primarily leveraged to map the positions of the corresponding RGB images within the environment, facilitating the construction of local or global grid-based representations. In contrast, our approach directly incorporates depth images as inputs, providing the agent with a clearer perception of the environmental layout. On the other hand, CONSOLE [30] leverages ChatGPT https://openai.com/blog/chatgpt to generate landmark cooccurrence priors, which are then aligned with the agent’s visual observations, while KERM [26] describes the attributes and relationships between objects as textual knowledge. Despite the varied forms of incorporating textual semantics, these methods typically compact textual semantics with corresponding visual views before feeding them to the agent. In contrast, our approach treats the generated textual semantics as a separate environmental representation, enabling the agent to better ground environments with textual instructions, thereby significantly improving both navigation success and object recognition accuracy. In summary, compared to previous approaches, we improve spatial-semantic perception for agent by equipping agent with more alternative environmental representations.
Figure 7: The weights for hybrid representations in the HRF module change during training (see Eq. 6).
Figure 8: Similarity matrix for the top 50 most frequent landmarks: REVERIE instructions vs. environments (left), R2R instructions vs. environments (right). Best viewed in color.
7.2 Fusion Weights of Hybrid Representations.
We quantitatively illustrate the weights (see Eq. 6) assigned to each branch of the HMF module during training (in Fig. 7) and inference (in Tab. 9). These weights are normalized between 0 and 1, with higher values denoting stronger dependency on the corresponding environmental representation. As shown in Fig. 7, initially, all four representations have approximately equal weights of 0.25 during training, reflecting random initialization and balanced contributions. Over the course of training, the global RGB representation weight $\beta_{4}$ increases, while those of the textual semantic $\beta_{1}$ and depth $\beta_{3}$ modalities decrease. This trend arises because textual semantics and depth provide sparse yet precise information, which is particularly beneficial in the early stages of training. Despite the increasing dominance of the RGB modality, the agent’s reliance on textual semantics and depth remains above 20%, underscoring their essential role in navigation. Tab. 9 also demonstrates that the agent effectively utilizes the introduced depth spatial and textual semantic information during inference, even though it exhibits a stronger reliance on the RGB modality.
Splits Local Global
Semantic $\beta_{1}$ RGB $\beta_{2}$ Depth $\beta_{3}$ RGB $\beta_{4}$
Validation Seen 0.2147 0.2458 0.2239 0.3156
Validation Unseen 0.2151 0.2662 0.2216 0.2971
Table 9: Variation in fusion weights (in Eq. 6) of hybrid representations during inference on R2R. $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ denotes the fusion weight of textual semantic, local RGB, depth, global RGB representations in the HRF module, respectively.
7.3 Quality of Generated Salient Landmarks
The quality of the extracted salient landmarks is crucial for the agent to accurately understand the textual semantics of the environment. In Fig. 8, we compute the similarity between the top-50 most frequently occurring landmarks in our generated salient landmarks and those from R2R [3] and REVERIE [40] to evaluate the quality of the generated salient landmarks. While the types and frequencies of these generated landmarks is different from those in the datasets, they can contain and convey most similar sematic (similarities surpass 0.5). Notably, the plug-and-play CLIP text encoder excels in understanding image-related descriptions, while BERT is more adept at comprehending natural language sentences. Therefore, we use the CLIP text encoder to extract features from the generated salient landmarks, instead of BERT.
Figure 9: Analysis of the contrastive loss weight $\lambda_{2}$ in HRF module on R2R validation unseen split.
7.4 Analysis of the Contrastive Loss Weight
Our ablation studies indicate that the contrastive learning loss weight, denoted as $\lambda_{2}$ in Eq. 11, significantly affects model navigation performance. As shown in Fig. 9, we observe that, within a certain range, increasing the contrastive loss weight improves navigation performance. However, when the contrastive loss weight is too high, performance (SR/OSR) declines, thus overemphasizing contrastive learning loss may lead the agent to prioritize modality alignment over core sequential navigation tasks.
ID DSP TSU HRF SR SPL RGSPL
1 46.98/52.51 33.73/36.06 23.03/22.06
2 52.00/52.53 39.21/39.53 26.54/24.97
3 53.17/53.89 38.31/40.89 26.73/26.43
4 55.07/51.97 39.50/37.87 26.70/25.20
5 51.75/54.39 38.86/41.54 26.56/27.31
Table 10: Ablation study on different components of SUSA on the REVERIE (validation/test) unseen split.
7.5 Overall Design on REVERIE
We further supplemented the performance of each key component of SUSA on the REVERIE task. To ensure fairness, we adopt the same hyperparameter setting ($\lambda_{2}$=0.8) as used in R2R, though this may be suboptimal for REVERIE. As shown in Tab. 10, our previous conclusions are reaffirmed: the DSP module allows the agent to better perceive spatial layouts, thus improving navigation efficiency (SPL). The TSU module enhances the agent’s understanding of textual semantics, resulting in higher success rates in both navigation (SR) and object recognition (RGS). An intriguing phenomenon, as shown in Tab. 10, is observed on the REVERIE dataset: directly incorporating the DSP and TSU modules (#4) results in an improvement on validation unseen (e.g., SR=55.07, SPL=39.50), while performance on test unseen deteriorates. In contrast, the addition of the HRF module (#5) yields opposite results. Overall, the introduction of the HRF module leads to superior navigation performance, both in terms of navigation (SR/SPL) and object grounding (RGSPL).
7.6 Computing Efficiency vs. Navigation Performance
Figure 10: Comparison of computational efficiency and navigation performance on R2R validation unseen split with several popular VLN methods.
As Fig. 10 illustrated, we compare the Giga Floating-point Operations (GFLOPs) and navigation performance (i.e., SR/SPL, which are more important metrics) of different VLN models, including DUET [5] , DSRG [53], HAMT [4] and our SUSA. To ensure a fair comparison across all methods, we performed single-step forward inference with a batch size of 8, an instruction length of 44, and 6 exploration map nodes. Compared to the baseline DUET, our approach introduces spatial and semantic environmental information, which inevitably increasing computational cost. However, despite this modest cost, the SUSA architecture notably enhances agent navigation performance.
8 Qualitative Examples
Failure cases. We provide an in-depth examination of failure cases, aiming to offer valuable insights for future model improvements. Fig. 11 displays the difference between the navigation trajectory predicted by the proposed SUSA and the ground truth trajectory. We found that, although textual semantics can provide the agent with a richer understanding, it may introduces ambiguity in certain scenarios. Concretely, when two identical salient landmarks appear in different scenes (e.g., the staircase and bathroom in Fig. 11), the agent may struggle to distinguish the correct navigable node, leading to erroneous navigation decisions.
More Qualitative Examples. We further visualize first-person navigation trajectories on the R2R validation unseen split, as illustrated in Fig. 12.
Figure 11: Top-down trajectories visualization of failed cases. The corresponding instructions are shown at the bottom. Partial views of final stopping position and corresponding landmark captions of ground truth and SUSA are also provided
9 Limitations and Future Work
As previously mentioned, when an agent encounters the same landmark at different navigable points, it may struggle to make confident navigation decisions, leading to failures. To address this limitation, 1) our future research will explore adjusting the confidence levels associated with different environmental representations to help the agent make more accurate and reliable predictions. 2) Beyond landmarks, we will construct direction-related textual semantics (e.g., “past,” “turn left”) to better model the spatial relationships between actions in instructions and the environmental layout or historical exploration maps. We also intend to investigate how well our approach generalizes to other navigation tasks. Additionally, given the complexity of VLN tasks, current mainstream VLN frameworks are often intricate and cumbersome, with a wide variety of evaluation metrics that are difficult to balance. Ongoing research in the future is expected to develop more lightweight VLN model architectures and comprehensive evaluation metrics.
Figure 12: First-person trajectory visualizations of DUET (left) and our SUSA (right), with the corresponding instructions provided below. Red arrows indicate the next selected viewpoint from current position.

### v4

@@
 Navigating unseen environments based on natural language instructions remains difficult for egocentric agents in Vision-and-Language Navigation (VLN).
-While recent advancements have yielded promising outcomes, they primarily rely on RGB images for environmental representation, often overlooking the underlying semantic knowledge and spatial cues.
-Intuitively, humans inherently ground textual semantics within the spatial layout during indoor navigation.
-Inspired by this, we propose a versatile Semantic Understanding and Spatial Awareness (SUSA) architecture to facilitate navigation.
-SUSA includes a Textual Semantic Understanding (TSU) module, which narrows the modality gap between instructions and environments by generating and associating the descriptions of environmental landmarks in the agent’s immediate surroundings.
-Additionally, a Depth-based Spatial Perception (DSP) module incrementally constructs a depth exploration map, enabling a more nuanced comprehension of environmental layouts.
-Experimental results demonstrate that SUSA’s hybrid semantic-spatial representations effectively enhance navigation performance, setting new state-of-the-art performance across three VLN benchmarks (REVERIE, R2R, and SOON).
-The source code are publicly available at https://github.com/HCI-LMC/VLN-SUSA.
+Existing approaches primarily rely on RGB images for environmental representation, underutilizing latent textual semantic and spatial cues and leaving the modality gap between instructions and scarce environmental representations unresolved.
+Intuitively, humans inherently ground semantic knowledge within spatial layouts during indoor navigation.
+Inspired by this, we propose a versatile Semantic Understanding and Spatial Awareness (SUSA) architecture to encourage agents to ground environment from diverse perspectives.
+SUSA includes a Textual Semantic Understanding (TSU) module, which narrows the modality gap between instructions and environments by generating and associating the descriptions of environmental landmarks in agent’s immediate surroundings.
+Additionally, a Depth-enhanced Spatial Perception (DSP) module incrementally constructs a depth exploration map, enabling a more nuanced comprehension of environmental layouts.
+Experiments demonstrate that SUSA’s hybrid semantic-spatial representations effectively enhance navigation performance, setting new state-of-the-art performance across three VLN benchmarks (REVERIE, R2R, and SOON).
+The source code will be publicly available.
 1 Introduction
-With the advancement of multimodal technology [43, 32, 63, 47, 24], embodied VLN tasks [3, 40, 35, 37, 7, 44] have garnered significant attention due to their promising applications in robotics and intelligent assistance.
-Among them, conventional VLN tasks [3, 22, 20] focused on step-by-step navigation within unseen environments based on detailed instructions, while goal-oriented VLN tasks [65, 40, 45] demand agents to identify predefined objects based on high-level instructions.
-Consequently, agents must profoundly cognize and exploit the rich environmental information and ground textual instructions to brilliantly accomplish these arduous navigation tasks.
+With the advancement of multimodal technology [44, 47, 29, 59], several embodied VLN tasks [2, 41, 37, 8] have garnered significant attention due to their promising applications in robotics and intelligent assistance.
+Among them, conventional VLN tasks [2, 27] focused on step-by-step navigation within unseen environments based on detailed instructions, while goal-oriented VLN tasks [64, 41, 45] demand agents to identify predefined objects based on high-level instructions.
+Consequently, VLN agents are required to profoundly cognize the spatial information of environment and ground them with textual instructions to brilliantly accomplish arduous navigation tasks.
 Figure 1: An overview of the proposed SUSA.
 Unlike previous approaches that solely rely on RGB input, we introduce view-level textual panoramas and trajectory-level depth exploration maps, supporting the agent’s explicit understanding of the environment.
-A significant body of VLN research has focused on Transformer-based models [4, 18, 38, 55, 39] for interacting with instructions, environmental images, and historical trajectories.
-While these models have undergone substantial development and exhibit efficacy, two key challenges remain in VLN tasks.
-First, VLN agents struggle to precisely match the landmarks (e.g., “fireplace” in Fig.
-1) in instructions with the entities in environments, due to the intrinsic modality heterogeneity between vision and language.
-To address this, recent work [11, 54, 28] has leveraged contrastive learning to promote alignment, while others have developed repositories of candidate object concepts [53, 29] or utilized large language models (LLMs) to identify the most relevant landmarks in the visual scene [30].
-However, these methods often lack an explicit understanding of environments from the perspective of textual semantics (i.e., the linguistic descriptions of landmarks).
-Intuitively, when navigating indoors or searching for objects, humans naturally associate the textual semantics of their surroundings.
-The second challenge pertains to the scarcity of diverse environmental sources, which often leads VLN agents to overfit to familiar environments, hindering generalization.
-To alleviate this limitation, various methods [23, 57, 17, 6, 50, 27] have employed data augmentation to increase the environmental quantity and diversity.
-More recent endeavors enhance spatial perception by constructing finer-grained graph [25, 2], semantic map [15, 19], grid [58, 33], or volumetric [34] representations.
-However, these methods predominantly rely on RGB images, overlooking the depth images, which provides more intuitive spatial structure, especially in stereoscopic navigation environments.
-Furthermore, few approaches [48] treat depth as a distinct modality capable of offering complementary spatial information, instead using it merely as an auxiliary signal to refine RGB-based representations.
-In conclusion, the co-occurrence semantic and spatial environment representations, beyond surface RGB images, has not been fully explored or exploited.
-In this work, we propose the Semantic Understanding and Spatial Awareness (SUSA) architecture, designed to explicitly capture both semantic and spatial information during navigation.
+Considerable literature has exhibited the efficacy of Transformer-based models [5, 20, 38, 52, 40] for interacting with instructions, RGB-based environmental observations, and historical trajectories.
+However, we suggest that RGB-only visual environment representations leave VLN fraught with two challenges: the modality gap between instructions and the environment, and the scarcity of training environmental representations.
+Firstly, VLN agents struggle to precisely match the landmarks in language instructions with the entities in visual environments, due to the intrinsic modality heterogeneity.
+To address this, recent works [12, 32] utilized contrastive learning to promote alignment, while others leveraged candidate object concepts [50, 33], large language models (LLMs) [34] or retrieved knowledge [31] to explicitly reason the textual semantics (i.e., linguistic descriptions about landmarks) to enhance environmental representations.
+However, these methods do not directly align textual semantics with landmarks (e.g., “fireplace” in Fig.
+1) in instruction, yielding trivial gain.
+The second challenge pertains to the monotonous and limited environmental representations, which often leads agents to overfit to familiar spatial layouts [51].
+To alleviate this limitation, recent endeavors [55, 7] augment the quantity of environmental scenes, while others diversify environmental representations by constructing graphs [16, 23], grids [56, 1], or volumetric environments [36].
+However, these RGB-based methods easily suffer from attention imbalance or environmental bias [65, 61].
+FDA [18] utilizes frequency information to enrich environmental representations.
+However, depth images vividly convey spatial information (e.g., doors and corridors), which is often overlooked in discrete navigation tasks [49].
+Few works [1, 56] incorporate depth, they merely project depth values onto the corresponding spatial locations of RGB images, rather than independently grounding depth[53].
+Overall, the underlying textual semantic and spatial representations remains underutilized, though studies [22] suggest that injecting inductive biases where different modules access distinct modalities, offers a promising solution.
+In this work, we propose the Semantic Understanding and Spatial Awareness (SUSA) architecture, which individually grounds semantic and spatial representations of the environment with instructions and synergistically predicts navigation actions from diverse perspectives.
 As illustrated in Fig.
-1, we first construct a textual semantic panorama by extracting salient landmarks as the textual semantics for each view to explicitly match the textual semantics with language instructions, and select the most instruction-relevant view for navigation prediction, thereby bridging the semantic gap between the visual and linguistic modalities.
-To enhance the spatial awareness, we then progressively construct an independent depth exploration map, which helps mitigate the risk of the agent overfitting to the information-dense RGB images.
-To enabling the agent to absorb complementary information from different environmental perspectives, we adaptively fuse these hybrid semantic-spatial environmental representations to facilitate informed navigation predictions, and employ contrastive learning to align these hybrid representations with language instructions.
+1, we first construct a view-level textual panorama by extracting salient landmarks as the textual semantics to explicitly match them with language instructions, and select the most instruction-relevant view, thereby bridging the semantic gap between the visual and linguistic modalities.
+Further, beyond RGB images, we progressively construct an independent trajectory-level depth exploration map to enhance the agent’s spatial perception of the environment, also directly aligning it with instructions.
+To enable the agent to absorb complementary information from different environmental perspectives, we aggregate these hybrid semantic-spatial environmental representations to synergistically facilitate hierarchical navigation predictions, and employ contrastive learning to align these hybrid representations with language instructions.
 We empirically substantiate that exclude semantic panoramas and depth exploration maps during pretraining manifest markedly enhanced generalization capabilities.
@@
 The main contributions of our SUSA are as follows:
-  • SUSA devises a textual-aware semantic understanding module to match the most related landmarks between environments and instructions, explicitly addressing the modality heterogeneity.
-  • SUSA presents a depth-based spatial perception module to explore and represent environmental layouts beyond surface RGB content.
+  • SUSA devises a textual-aware semantic understanding (TSU) module to match the most related landmarks between environments and instructions, explicitly addressing the modality heterogeneity.
+  • SUSA further presents a depth-enhanced spatial perception (DSP) module to explore and represent environmental layouts beyond surface RGB content.
   • Extensive experiments on three VLN benchmarks have demonstrated the superiority of hybrid semantic-spatial environmental representations introduced by SUSA.
@@
 Vision-and-Language Navigation (VLN).
-Vision-and-Language Navigation (VLN) tasks [3, 40, 65] require an agent to navigate sequentially towards a designated location or object based on natural language instructions.
-Given that the VLN task can be formulated as a partially observable Markov decision process, early methods [56, 52, 66, 13, 21, 36] primarily relied on reinforcement and imitation learning paradigms to improve navigation performance, incorporating various strategies and auxiliary loss [64, 56] designs.
+Vision-and-Language Navigation (VLN) tasks [2, 41, 64] require an agent to navigate in embodied environments sequentially towards a designated location or object based on natural language instructions.
+Given that the VLN task can be formulated as a partially observable Markov decision process, early methods [54, 66, 14, 26] primarily relied on reinforcement learning and imitation learning paradigms to improve navigation performance, incorporating various strategies and auxiliary loss [63, 54] designs.
 More recently, to better generalize to unknown environments and ground instructions, many researchers [5, 20, 38, 52, 40] have focused on transformer-based architectures, which effectively model interactions among instructions, environmental features, and historical navigation trajectories.
-While these approaches have propelled VLN forward, the co-occurrence environmental sources have not been fully utilized.
-This paper introduces text semantics and depth information to enrich environmental representations.
+While these approaches have propelled VLN forward, they mainly rely on fragile RGB-based environment representations.
+This paper encourages the agent to perceive and ground the environment from text semantics and depth perspectives.
 Semantic Alignment in VLN.
-To align the relationship between the environment and instructions, current researchers and practitioners in the VLN community can be divided into two main categories: implicit and explicit approaches.
-Implicit methods [11, 54, 28] commonly employ contrastive learning to associate visual contexts with textual instructions or use attention mechanisms to model the relationship.
-However, due to the inherent semantic gap between these modalities, implicit approaches have limited alignment capabilities.
-Explicit alignment techniques aim to project both semantics into a unified modality space.
-For example, some methods transform the environment into a text-based modality, Console [30] scores landmark co-occurrences to identify key navigation objects, while AACL [29] leverages CLIP [43] to refine textual entities, compacting them with visual features.
-However, these explicit methods often directly “stitch” textual and visual environment representations, which may leading to an overreliance on precise but abstract text while underutilizing rich yet redundant visual data.
-We thus use environmental textual semantic as an independent modality to explicitly match the most instruction-relevant landmarks.
-Environment Representation in VLN.
-Limited training data and embodied environment scarcity in VLN tasks usually hinder the agent to generalize the unknown scenes or layouts.
-Previous approaches mitigate this issue through data augmentation [31, 57] or by enhancing visual representations[61, 34].
-ScaleVLN [57] introduces thousands of photorealistic environments and FDA [17] incorporates low-frequency interference images to emphasize essential high-frequency details.
-Advanced visual representations, such as topological maps and grid-based layouts, further enable agents to retain temporal navigation paths and more precisely perceive spatial structures [1, 58].
-For example, DUET [5] employs dual-scale transformers [51] to balance global and local decision-making, while VER [34] adopts volumetric models to represent 3D spatial structures.
-While most research relies on RGB representations, which tend to capture redundant appearance features, our method incorporates depth data, offering agents a more intuitive and spatially aware understanding of the environment.
+To align the semantic relationship between the environment and instructions, current practitioners in the VLN community can be divided into two categories: implicit and explicit approaches.
+Implicit methods [12, 51, 32] commonly employ contrastive learning or attention mechanisms to associate visual contexts with textual instructions.
+However, given the inherent semantic gap between these modalities, implicit approaches have limited alignment capabilities.
+Explicit alignment techniques aim to project both semantics into a same modality space.
+For example, some methods transform the environment into a text-based modality, Console [34] scores landmark co-occurrences to identify key objects, and AACL [33] leverages CLIP [44] to refine textual entities.
+Similar to our approach, KERM [31] utilizes retrieved knowledge as textual semantics for each view.
+However, these explicit methods directly “stitch” textual and visual environment representations, which may lead to an overreliance on text modality yet underutilizing visual data [67].
+We thus take environmental textual semantic as an independent modality to directly match the most instruction-relevant landmarks.
+Environment Representations in VLN.
+Limited training sources and environmental representations in VLN tasks hinder the agent to generalize the unknown layouts.
+Prior endeavors mitigate this limitation through data augmentation [35, 55] or by diversifying visual representations [60, 1, 56, 36].
+However, these approaches primarily rely on RGB images, while alternative modalities such as frequency [18], spectral[24], depth [49] and semantic information [53, 21, 4] remain relatively underexplored.
+Among them, SEAT [53] enhances cross-domain environmental perception by querying the Matterport3D simulator[3] for depth and semantics.
+However, SEAT fuses these multi-type representations into a unified space without explicitly grounding each modality.
+Moreover, these semantic representations [53, 21] are not inherently textual, limiting their ability to bridge the modality gap.
+In contrast, our method constructs the view-level textual panoramas and trajectory-level depth exploration maps, directly grounding them with instructions to perceive rich environmental representations.
 Figure 2: The detailed architecture of our proposed SUSA model.
 The orange and green arrows highlight our proposed Textual-Aware Semantic Understanding (TSU) and Depth-Aware Spatial Perception (DSP) modules, respectively.
-The Hybrid Representation Fusion (HRF) module is designed for aligning environmental representations with instructions.
+The Hybrid Representation Aggregation (HRA) module is designed for aligning environmental representations with instructions.
 3 Methodology
@@
 Task Formulation.
-We focus on discrete indoor VLN tasks [3, 40, 65] within a Matterport3D simulator [3], modeled as an undirected graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}$ denotes nodes and $\mathcal{E}$ signifies the connectivity paths.
-At the outset of each episode, the agent, equipped with a GPS sensor, RGB and depth cameras, is situated at a randomly navigable node and receives a language instruction $I=\{\text{w}_{i}\}_{i=1}^{l}$ including $l$ words.
+We focus on discrete VLN tasks [2, 41, 64] within a Matterport3D simulator [3], modeled as an undirected graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}$ denotes nodes and $\mathcal{E}$ signifies the connectivity paths.
+At the outset of each episode, the agent equipped with a GPS sensor, RGB and depth cameras, is situated at a randomly navigable node and receives a language instruction $I=\{\text{w}_{i}\}_{i=1}^{l}$ including $l$ words.
 We use BERT [10] to extract instruction features and then embed them as =$\mathcal{I}=\{w_{1},\dotso,w_{l}\}$.
 At each time step $t$, the agent can perceive the surrounding RGB panorama $R_{t}=\{\text{r}_{t,i}\}_{i=1}^{n=36}$ and depth panorama $D_{t}=\{\text{d}_{t,i}\}_{i=1}^{n=36}$ along with $n$ corresponding views.
-During navigation, the agent must decide its next action $a_{t}$ by selecting one navigable node from the candidate actions $\mathcal{A}_{t}=\{a_{t,i}\}_{i=0}^{k}$.
+During navigation, the agent make the next action $a_{t}$ by selecting one navigable node from the candidate actions $\mathcal{A}_{t}=\{a_{t,i}\}_{i=0}^{k}$.
 The sequential decision-making process is iterated until the agent decides to stop at the current location or exceeds the maximum navigation steps.
@@
 Overview of the Proposed Method.
-Our objective is to equip the agent with a textual semantic panorama and a depth exploration map to accord complementary environmental content.
-We adopt DUET [5] as our baseline, which merely utilizes RGB panoramas and topological maps as environmental representations for both local and global action predictions.
+Our objective is to accord complementary environmental content by equipping the agent with a textual semantic panorama and a depth exploration map.
+Following prevailing approaches [18, 30, 19], we adopt DUET [6] as our baseline, which merely utilizes RGB panoramas and topological maps as environmental representations for both local and global action predictions.
 The SUSA framework consists of three main modules: (a) textual-aware semantic selection is detailed in Sec.
 3.2, (b) depth-aware visual environment exploration is presented in Sec.
-3.3, and (c) hybrid representation fusion and action prediction is subsequently introduced in Sec.
+3.3, and (c) hybrid representation aggregation and hierarchical action prediction is subsequently introduced in Sec.
 3.4, as illustrated in Fig.
@@
 Additionally, we employ the CLIP (ViT-L-14-336px) text encoder [44] to extract textual semantic panorama features, embedding them as $\mathcal{X}^{\mathcal{S}}_{t}=\{x^{s}_{t,1},\dots,x^{s}_{t,n}\}$.
-During navigation, we leverage the frozen BLIP-2 and CLIP models to empower the agent to obtain real-time textual semantic features from the navigable nodes in its vicinity.
+Thanks to the powerful yet frozen BLIP-2 and CLIP models, we empower the agent to obtain textual semantic features from the navigable nodes in its vicinity during navigation.
 Textual Semantic Selector.
-Previous methods [53, 29] aligned environmental RGB images with instructions by exploiting textual semantics yet the inherent modality discrepancy between visual and language modalities remained unaddressed.
-In contrast, we treat the textual semantic as a distinct modality, directly matching salient landmarks with the instructions.
+Previous methods [50, 33, 31] aligned environmental RGB images with instructions by exploiting textual semantics yet the inherent modality discrepancy between visual and language modalities remained unaddressed.
+Similar to [39], we treat the textual semantic as a perceptual representation for navigation in this module, directly matching salient landmarks with the instructions.
 As shown in Fig.
@@
 Specifically, we utilize a standard multi-layer transformer decoder to dynamically match the relevance $\Gamma_{t}^{dyn}$ between textual semantics and instructions, as formalized by:
-$\small\begin{split}\Gamma_{t}^{dyn}=\text{TCA}(\mathcal{S}_{t},\mathcal{I},\mathcal{I}),\end{split}$ (2)
+$\small\begin{split}\Gamma_{t}^{dyn}=\text{TCA}(\mathcal{X}^{\mathcal{S}}_{t},\mathcal{I},\mathcal{I}),\end{split}$ (2)
 where TCA denotes a textual modality cross-attention mechanism that models textual semantic-instruction relationships.
@@
 where the balance factor $\delta$ regulates the relative contributions of each matching strategy.
-By mapping the environment to the same textual modality, the textual semantic selector enables the agent to better understand the environment context from a textual semantic perspective, thus facilitating grounded language instructions.
-3.3 Depth-Aware Spatial Perception
-Depth images, which also contain more intuitive and easily distinguishable structural information, were not adequately exploited.
-To enhance the spatial perception for agent, we concurrently construct a depth-based exploration map to memorize depth-based navigation trajectory alongside the RGB-based one [5].
-Depth-Based Exploration Mapping.
+By mapping the environment to the same textual modality, the textual semantic selector enables the agent to explicitly understand the environment context from a textual semantic perspective, thus facilitating grounded language instructions.
+3.3 Depth-Enhanced Spatial Perception
+Depth images, which also contain more intuitive and easily distinguishable structural information, were not adequately exploited in earlier works.
+To enhance the spatial perception, we concurrently construct a depth-based exploration map to memorize depth-based navigation trajectory alongside the RGB-based one [6].
+Depth-Enhanced Exploration Mapping.
 Concretely, we obtain ground truth depth images for each view from the Matterport3D simulator.
 As illustrated in Fig.
-2, we then employ CLIP (ViT-L/14@336px) and ResNet-50 [16] (pretrained on the Gibson [59]) as the depth and RGB encoders, respectively, to extract panoramic depth features $\mathcal{D}_{t}\in\mathbb{R}^{n\times d_{v}}$ and panoramic RGB features $\mathcal{R}_{t}\in\mathbb{R}^{n\times d_{v}}$, where $d_{v}$ represents the dimension of each view.
+2, we then employ ResNet-50 [17] (pretrained on the Gibson [57]) and CLIP (ViT-L/14@336px) as the depth and RGB encoders, respectively, to extract panoramic depth features $\mathcal{D}_{t}\in\mathbb{R}^{n\times d_{v}}$ and panoramic RGB features $\mathcal{R}_{t}\in\mathbb{R}^{n\times d_{v}}$, where $d_{v}$ represents the dimension of each view.
 Moreover, we employ a two-layer transformer-based panorama encoding module to perform self-attention and sequentially encode the depth and RGB panoramic images.
@@
 Subsequently, we aggregate each visited node and its adjacent views by average pooling to construct depth and RGB exploration maps in parallel.
-The explored depth and RGB trajectory maps are formalized as $\mathcal{T}^{\mathcal{D}}\in\mathbb{R}^{k\times d_{n}}$ and $\mathcal{T}^{\mathcal{R}}_{t}\in\mathbb{R}^{k\times d_{n}}_{t}$ Here, $k$ denotes the number of navigable nodes at step $t$, while $d_{n}$ represents the dimension of each navigable node.
-These exploration maps memorize the agent navigation trajectory information for global action predictions, which serves to guide the agent traversal across various navigable points when considering potential backtracking.
+The explored depth and RGB trajectory maps are formalized as $\mathcal{T}_{t}^{\mathcal{D}}\in\mathbb{R}^{k\times d_{n}}$ and $\mathcal{T}^{\mathcal{R}}_{t}\in\mathbb{R}^{k\times d_{n}}_{t}$ Here, $k$ denotes the number of navigable nodes at step $t$, while $d_{n}$ represents the dimension of each navigable node.
+These exploration maps memorize the agent navigation trajectory information for global action predictions, which serves to guide the agent traversal across various navigable nodes when considering backtracking.
 Cross-Modal Interaction and Reasoning.
 To defer to instructions, we utilize a multi-layer cross-encoder that interacts with both the language instructions and the depth exploration map features to generate the refined depth map $\tilde{\mathcal{T}^{\mathcal{D}}}$.
-This encoder comprises four residual connection layers of a cross-modal transformer, each incorporating graph-aware self-attention [5] (GASA), cross-modal attention, and a feed-forward network.
-The GASA operation is formally expressed as $\text{GASA}(X)=\eth(X\Theta_{q}(X\Theta_{k})^{T}+\mathcal{M})X\Theta_{v}$ where $\mathcal{M}$ represents a learnable distance matrix encoding the relative distances between navigation points, $\eth$ denotes the Softmax activation function, and $\Theta_{q},\Theta_{k},\Theta_{v}$ are learnable weight matrices.
+This encoder comprises four residual connection layers of a cross-modal transformer, each incorporating graph-aware self-attention (GASA) [6], cross-modal attention, and a feed-forward network.
+The GASA operation is formally expressed as $\text{GASA}(X)=\eth(X\Theta_{q}(X\Theta_{k})^{T}+\mathcal{M})X\Theta_{v}$ where $\mathcal{M}$ represents a learnable distance matrix encoding the relative distances between navigation nodes, $\eth$ denotes the Softmax activation function, and $\Theta_{q},\Theta_{k},\Theta_{v}$ are learnable weight matrices.
 In parallel, we facilitate similar cross-modal interactions between the language instructions and the RGB exploration map $\mathcal{T}^{\mathcal{R}}$, as well as the RGB panoramas $\mathcal{X}^{\mathcal{R}}$.
 These are encoded as $\tilde{\mathcal{T}^{\mathcal{R}}}$ and $\tilde{\mathcal{X}^{\mathcal{R}}}$, thereby enabling the model to perform both global and local predictions based on the given instructions $\mathcal{I}$.
-Compared to the standalone $\mathcal{T}^{\mathcal{R}}$, the incorporation of a depth exploration map $\mathcal{T}^{\mathcal{D}}$ enhances the perceptual clarity of environmental structures, mitigating the risk of overfitting to RGB-based visual noises and biases.
-Figure 4: Illustration of the proposed HRF module in  Fig. 2.
-3.4 Hybrid Representation Fusion and Prediction
+Compared to the standalone $\mathcal{T}^{\mathcal{R}}$, the incorporation of a depth exploration map $\mathcal{T}^{\mathcal{D}}$ enhances the perceptual clarity of environmental structures, mitigating the risk of overfitting to RGB-based visual noises [25] or biases [22, 51].
+Figure 4: Illustration of the pipeline for fusing hybrid environmental representations in the proposed HRA module in Fig. 2.
+3.4 Hybrid Representation Aggregation
 Building on aforementioned methods, at the current state $t$, the agent synchronously perceives the environment through multiple modalities—textual semantic, depth, and RGB —at various scales (local and global).
@@
 Specifically, the attention pooling AP operation is defined as follows:
-^$\hat{\mathcal{H}}=\text{AP}(\mathcal{H},v),\quad\mathcal{H}\in\{\mathcal{X}^{\mathcal{S}},\mathcal{T}^{\mathcal{D}},\mathcal{X}^{R},\mathcal{T}^{R},\mathcal{I}\},$ (4)
+^$\hat{\mathcal{H}}=\text{AP}(\mathcal{H},v),\quad\mathcal{H}\in\{\tilde{\mathcal{X}^{\mathcal{S}}},\tilde{\mathcal{T}^{\mathcal{D}}},\tilde{\mathcal{X}^{R}},\tilde{\mathcal{T}^{R}},\mathcal{I}\},$ (4)
 where $v\in\mathbb{R}^{d}$ denoting the learnable query vector and ^$\hat{\mathcal{H}}\in\mathbb{R}^{d}$. The AP operator is given by:
 $\eta=\eth\left(\tau(\mathcal{H})\otimes v^{\top}\right),\quad\hat{\mathcal{H}}=\tau\left(\eta\odot\mathcal{H}\right),$ (5)
-with $\tau$ representing the tanh activation function. Subsequently, we perform weighted fusion of the hybrid environmental representations $E^{\mathrm{hyb}}$, as follows:
-$\small\footnotesize\mathcal{B}=\begin{bmatrix}\beta_{1}&\beta_{2}&\beta_{3}&\beta_{4}\end{bmatrix}^{\top}=\sigma(\mathrm{FFN}[\mathcal{X}_{0}^{\mathcal{S}};\mathcal{X}_{0}^{\mathcal{R}};\mathcal{T}_{0}^{\mathcal{D}};\mathcal{T}_{0}^{\mathcal{R}}])$ (6)
-$\small\footnotesize E^{\mathrm{hyb}}=\begin{bmatrix}\hat{\mathcal{X}^{\mathcal{S}}}&\hat{\mathcal{X}^{\mathcal{R}}}&\hat{\mathcal{T}^{\mathcal{D}}}&\hat{\mathcal{T}^{\mathcal{R}}}\end{bmatrix}\cdot\mathcal{B}$ (7)
+with $\tau$ representing the tanh activation function. Subsequently, we perform weighted aggregation of the hybrid environmental representations $E^{\mathrm{hyb}}$, as follows:
+$\small\footnotesize\mathcal{B}=\begin{bmatrix}\beta_{1}&\beta_{2}&\beta_{3}&\beta_{4}\end{bmatrix}^{\top}=\sigma(\mathrm{FFN}[\tilde{\mathcal{X}_{0}^{\mathcal{S}}};\tilde{\mathcal{X}_{0}^{\mathcal{R}}};\tilde{\mathcal{T}_{0}^{\mathcal{D}}};\tilde{\mathcal{T}_{0}^{\mathcal{R}}}])$ (6)
+$\small\footnotesize E^{\mathrm{hyb}}=\begin{bmatrix}\hat{\mathcal{X}^{\mathcal{S}}}&\hat{\mathcal{X}^{\mathcal{R}}}&\hat{\mathcal{T}^{\mathcal{D}}}&\hat{\mathcal{T}^{\mathcal{R}}}\end{bmatrix}\cdot\mathcal{B},$ (7)
 where $\sigma$ stands for the sigmoid activation function, and [;][;][ ; ] denotes concatenation.
@@
 Object Grounding (OG) task to localize target objects in REVERIE and SOON.
-Given the sparsity of depth images and textual semantics compared to RGB images, completely pre-train the model may lead to overfitting and increased training costs.
+Given the sparsity of depth images and textual semantics compared to RGB images, completely pre-train all parameter of SUSA may lead to overfitting and increased training costs.
 As shown in Fig.
@@
 HOP+ [43] 10.59 55.87 49.55 40.76 36.22 14.57 36.07 31.13 22.49 19.33 15.17 33.82 28.24 20.20 16.86
-DUET [5] 13.86 71.75 63.94 57.41 51.14 22.11 46.98 33.73 32.15 23.03 21.30 52.51 36.06 31.88 22.06
 DSRG [50] - 75.69 68.09 61.07 54.72 - 47.83 34.02 32.69 23.37 - 54.04 37.09 32.49 22.18
@@
 VER [36] 16.13 75.83 66.19 61.71 56.20 23.03 55.98 39.66 33.71 23.70 24.74 56.82 38.76 33.88 23.19
+DUET [6] 13.86 71.75 63.94 57.41 51.14 22.11 46.98 33.73 32.15 23.03 21.30 52.51 36.06 31.88 22.06
 SUSA (Ours) 14.60 76.95 69.07 61.77 55.86 22.59 51.75 38.86 35.02 26.56 17.86 54.39 41.54 36.11 27.31
@@
 During fine-tuning, we employ a contrastive learning loss, $\mathcal{L}_{cl}$ to align the hybrid environmental representations $E^{hyb}$ with the corresponding instructions $\mathcal{I}$:
-$\textstyle\mathcal{L}_{cl}=-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{j})/t)}{\sum_{j=1}^{B}\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{j})/t)}$ (10)
-$\textstyle-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((\mathcal{I}_{i},E^{\mathrm{hyb}}_{j})/t)}{\sum_{j=1}^{B}\exp((\mathcal{I}_{i},E^{\mathrm{hyb}}_{j})/t)},$
+$\textstyle\mathcal{L}_{cl}=-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{i})/t)}{\sum_{j=1}^{B}\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{j})/t)}$ (10)
+$\textstyle-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((\mathcal{I}_{i},E^{\mathrm{hyb}}_{i})/t)}{\sum_{j=1}^{B}\exp((\mathcal{I}_{i},E^{\mathrm{hyb}}_{j})/t)},$
 where $B$ and $t$ represent the batch size and the temperature.
 Then, we train the model based on hybrid environment representations by performing single-step action prediction [5, 6] with a contrastive learning loss $\mathcal{L}_{cl}$.
-The hybrid single action prediction loss in behavior cloning is computed as $\mathcal{L}_{hsap}^{*}=\sum_{t=1}^{T}-\log p(a_{t}^{*}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$ and $\mathcal{L}_{hsap}^{\star}=\sum_{t=1}^{T}-\log p(a_{t}^{\star}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$, where $a_{t}^{*}$ represents the ground truth action, and $a_{t}^{\star}$ denotes the pseudo-target action sampled from current exploration map.
+The hybrid single action prediction loss in behavior cloning is computed as $\mathcal{L}_{hsap}^{gt}=\sum_{t=1}^{T}-\log p(a_{t}^{gt}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$ and $\mathcal{L}_{hsap}^{\star}=\sum_{t=1}^{T}-\log p(a_{t}^{\star}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$, where $a_{t}^{gt}$ represents the ground truth action, and $a_{t}^{\star}$ denotes the pseudo-target action sampled from current exploration map.
 The final training loss is then expressed as follows:
-$\mathcal{L}_{SUSA}=\lambda_{1}\mathcal{L}_{hsap}^{*}+\mathcal{L}_{hsap}^{\star}+\lambda_{2}\mathcal{L}_{cl}$ (11)
+$\mathcal{L}_{SUSA}=\lambda_{1}\mathcal{L}_{hsap}^{gt}+\mathcal{L}_{hsap}^{\star}+\lambda_{2}\mathcal{L}_{cl}$ (11)
 where $\lambda_{1}$ and $\lambda_{2}$ denote loss weights.
-During the inference phase, the agent is tasked with sequentially navigating to the predicted navigable points, following the shortest path algorithm [12], without making any abrupt “jumps” between points.
+During the inference phase, the agent is tasked with sequentially navigating to the predicted navigable nodes, following the shortest path algorithm [13], without making any abrupt “jumps” between nodes.
 4 Experiments
@@
 For the REVERIE and SOON tasks, which involve object identification, we also evaluate object grounding metrics: Remote Grounding Success (RGS), and RGS weighted by Path Length (RGSPL).
+Given that SPL optimally balances navigation success rate and trajectory length, we employ it as our primary metric.
 Implementation Details.
@@
 The node and view dimensions are identical: $d_{v}=d_{n}=d=768$.
+We minimized changes to the baseline (DUET) [6] settings and used no additional annotations.
 All experiments were conducted on a single NVIDIA RTX 4090 GPU.
-We minimized changes to the baseline [5] settings and used no additional annotations.
 Detailed experimental setup is provided in the supplementary material.
@@
 We conduct a comprehensive comparison on the REVERIE [41], R2R [2], and SOON [64] datasets.
-For fairness, we exclude methods that involve pre-exploration [66, 46] or utilize large-scale data augmentation [6].
+For fairness, we exclude methods that involve pre-exploration [66, 46] or utilize large-scale data augmentation [7, 55].
 REVERIE.
@@
 2.
-Compared with the SOTA model GOAT [54] on the validation and test unseen splits, SUSA achieves 64.85% and 63.83% on SPL, respectively.
-These gains manifest not only an improved success rate but also enhanced navigation efficiency, since SPL considers both success rate and trajectory length.
+Compared with the SOTA ESCENE [62] model on the validation and test unseen splits, SUSA achieves 64.85% and 63.83% on SPL, respectively.
+These gains manifest not only an improved navigation performance but also enhanced navigation efficiency, since SPL considers both success rate and trajectory length.
 Methods Validation Unseen Test Unseen
@@
 HOP [42] 12.27 64 57 3.80 12.68 64 59 3.83
-DUET [5] 13.41 71 61 3.30 14.73 69 59 3.65
 DSRG [50] - 73 62 3.00 - 72 61 3.33
-AZHP [60] 13.68 71 60 3.25 14.47 69 59 3.43
 PanoGen [28] 13.40 74 64 3.03 14.38 71 61 3.31
@@
 FDA [18] 13.68 72 64 3.41 14.76 69 62 3.41
+KERM [31] 13.54 71.95 60.91 3.22 14.60 69.73 59.25 3.61
+AZHP [58] 13.68 71 60 3.25 14.47 69 59 3.43
 CONSOLE [34] 13.59 73 63 3.00 14.31 72 61 3.30
-KERM [26] 13.54 71.95 60.91 3.22 14.60 69.73 59.25 3.61
 ESCENE [62] 10.80 68 64 3.39 11.89 66 63 3.77
-GOAT* [54] 13.52 73.95 63.36 2.82 14.91 72.83 62.42 3.29
+DUET [6] 13.94 72 60 3.31 14.73 69 59 3.65
 SUSA (Ours) 12.18 73.05 64.85 3.06 13.27 72.56 63.83 3.23
@@
 SOON.
-On the SOON dataset, as evidenced in  Tab.
-3, SUSA outperforms the SOTA model GOAT by 2.27% and 0.29% on the validation and test unseen splits, respectively, in terms of SPL.
-Overall, the outstanding experimental results show that textual semantics boost navigation success via salient landmark matching, while the depth exploration map is beneficial to generalizable navigation efficiency.
+On the SOON dataset, as shown in Tab.
+3, SUSA achieves state-of-the-art performance across multiple metrics, including SR, SPL, and RGSPL, attaining SPL scores of 30.82% and 25.47% on the validation and test unseen splits, respectively.
+Moreover, SUSA consistently enhances the baseline (DUET)’s navigation performance and generalization on three datasets, underscoring the effectiveness of integrating textual semantics and depth exploration maps.
 Method Validation Unseen Test Unseen
 […]
-GBE [65] 28.54 19.52 13.34 1.16 21.45 12.90 9.23 0.45
+GridMM [56] 38.92 37.46 24.81 3.91 46.20 36.27 21.25 4.15
+SEAT [53] - 36.87 24.87 3.91 - 35.89 22.55 4.47
+KERM [31] 35.83 38.05 23.16 4.40 - - - -
 DUET [6] 36.20 36.28 22.58 3.75 41.83 33.44 21.42 4.17
-GridMM [58] 38.92 37.46 24.81 3.91 46.20 36.27 21.25 4.15
-KERM [26] 35.83 38.05 23.16 4.40 - - - -
-GOAT[54] - 40.35 28.05 5.85 - 40.50 25.18 6.10
 SUSA(Ours) 36.86 42.89 30.82 6.48 36.13 36.87 25.47 5.93
@@
 4.3 Diagnostic Experiment
-1) Overall Design.
 […]
-1 13.41 3.10 79.86 71.43 60.63
-2 13.85 3.08 80.42 72.80 62.71
-3 13.78 3.06 81.65 72.92 63.16
-4 12.18 3.06 79.99 73.05 64.85
-5 14.58 3.08 82.08 73.95 62.84
-Table 4: Ablation study on different components of SUSA on the R2R validation unseen split.
-To thoroughly evaluate the efficacy of key components, we conduct ablation experiments on the R2R validation unseen split in the Tab.
-4.
-As shown in (#2, #3, #4), both the Textual-aware Semantic Understanding (TSU) and Depth-aware Spatial Perception (DSP) modules consistently outperform the baseline (#1) [5], underscoring the substantial performance gains achieved by incorporating textual panoramas and depth exploration maps over relying solely on RGB.
-The hybrid representation fusion (HRF) module improves the OSR/SR score from 73.05%/79.99% to 73.95%/82.08%, demonstrating that the proposed HRF enables more precise alignment of environment-instruction pairs through contrastive learning.
-Analysis of the HRF module and the computational efficiency of the entire model is presented in supplementary materials.
-[…]
-ViT No 51.07 46.98 33.73 32.15 23.03
-CLIP No 57.17 51.55 35.75 35.19 24.40
-CLIP ResNet(ImageNet) 54.42 50.18 34.63 34.39 23.66
-CLIP ResNet(Gibson) 56.72 52.00 39.21 35.05 26.54
-Table 5: Ablation of different visual representations. Gibson and ImageNet serve as pre-training sources for the depth encoder.
+ViT $w/o$ 51.07 46.98 33.73 32.15 23.03
+CLIP $w/o$ 57.17 51.55 35.75 35.19 24.40
+CLIP ResNet-50(ImageNet) 54.42 50.18 34.63 34.39 23.66
+CLIP ResNet-50(Gibson) 56.72 52.00 39.21 35.05 26.54
+Table 4: Ablation of different visual representations on the REVERIE validation unseen split. Gibson and ImageNet serve as pre-training sources for the depth encoder in the DSP module.
 $\delta$ REVERIE R2R
@@
 Table 5: Ablation on the relative importance of static and dynamic matching in the TSU module, balanced by the balance factor $\delta$.
-2) Image Features vs.
+1) Image Features vs.
 Structural Features.
-We analyze various visual features (depth and RGB) contributions within the DSP module (Tab.
-5).
-For RGB images, substituting ViT [10] features with those from CLIP [43] yields marked improvements across all metrics, underscoring that CLIP achieves a better understanding of the environment than its ViT counterpart.
-As for depth features, we leveraged depth features extracted by ResNet-50 [16] pre-trained on the ImageNet [8] or Gibson [59] datasets to investigated the effect of depth exploration map.
-Compared to Matterport3D, the Gibson simulator offers more detailed and realistic physical environments along with higher-resolution visual information.
-The results demonstrate that simply incorporating depth information does not necessarily lead to improved performance.
-In comparison, the ResNet-50 pre-trained on the Gibson dataset extracts structural features that refine spatial perception, yielding notable improvements in the SPL metric without compromising SR or RGS.
+We analyze RGB and depth features contributions within the DSP module (Tab.
+4).
+For RGB images, substituting ViT [11] features with those from CLIP [44] yields marked improvements across all metrics, highlighting that CLIP achieves a better understanding of the environment than its ViT-based baseline counterpart.
+As for depth features, we leveraged depth features extracted by ResNet-50 [17] pre-trained on the ImageNet [9] or Gibson [57] datasets to investigated the effect of depth exploration maps.
+Compared to ImageNet, the Gibson simulator provides spatial structural information that serves as valuable navigation priors.
+Therefore, in  Tab.
+4, the ResNet-50 pre-trained on the Gibson extracts structural features that refine spatial perception, yielding notable improvements in the SPL metric without compromising SR or RGS.
 These improvements indicate that better spatial perception enhances navigation efficiency.
-3) Static Matching vs.
+2) Static Matching vs.
 Dynamic Matching.
@@
 Tab.
-6 summarizes these results, where w/o indicates the exclusion of the TSU module, corresponding to the results in Tab.
-4 (#2) and last row in Tab.
-5.
+5 summarizes these results, where w/o indicates the exclusion of the TSU module, corresponding to the last row results in Tab.
+4.
 Furthermore, adaptive denotes that $\delta$ is a learnable parameter which adaptively adjusts the contribution of each matching strategy.
 Compared to the first row in Tab.
-6, overall, incorporating the TSU module (regardless of static or dynamic matching strategy) improves SR and RGS, albeit at the cost of SPL on REVERIE.
+5, overall, incorporating the TSU module (regardless of static or dynamic matching strategy) improves SR and RGS, albeit at the cost of SPL on REVERIE.
 In contrast, it enhances SPL on R2R.
-While the TSU module can achieve superior performance under the adaptive setting, it falls short when $\delta$ = 0.5.
+While the TSU module can achieve superior performance under the adaptive setting, its performance is inferior when $\delta$ = 0.5.
 Specifically, at $\delta$ = 0.5, the SPL metric reaches 39.50% and 64.85% on the two datasets, with the RGS, nDTW, and sDTW metrics also achieving optimal performance.
 Excellent results demonstrate dual merits of proposed TSU module, which boosts object identification accuracy on the REVERIE task while simultaneously improving the fidelity between navigation trajectories and instructions on the R2R.
-Figure 5: a) and b) show the performance gap ($\downarrow$) between seen and unseen environments, while c) and d) present key metrics for different pretraining strategies on the R2R and REVERIE.
-4) Improved Generalization via Pretraining.
-As analyzed before, pretraining the entire SUSA framework on multiple auxiliary tasks may lead to overfitting, which can negatively impact the generalization in unseen environments, particularly considering the modality differences between textual panorama, depth exploration maps, and RGB representations.
+Figure 5: a) and b) show the performance gap ($\downarrow$) between seen and unseen environments, while c) and d) present key metrics ($\uparrow$) for different pretraining strategies on the R2R and REVERIE.
+3) Improved Generalization via Pretraining.
+As aforementioned in Sec.
+3.5, pretraining the entire SUSA framework on multiple auxiliary tasks may lead to overfitting, which may negatively impact the generalization in unseen environments, particularly considering the modality differences between textual panorama, depth exploration maps, and RGB representations.
 As shown in Fig.
-5, we observe that partial pretraining strategies not only consistently narrow the performance gap between seen and unseen environments but also outperform the completely/random pretraining strategies in navigation performance.
+5, we observe that the partial pretraining strategy not only consistently narrow the performance gap between seen and unseen environments but also outperform complete/random pretraining strategies in navigation performance.
 For example, as depicted in Fig.
-5 (a), on the REVERIE, the partial pretraining model achieves an SR gap of only 21.11% between the seen and unseen splits, reducing the gap by 5.92% compared to the complete pretraining strategy.
+5 (a), the partial pretraining model achieves an SR gap of only 21.11% between the seen and unseen splits on REVERIE, reducing the gap by 5.92% compared to the complete pretraining strategy.
 Additionally, as shown in Fig.
@@
 Fig.
-5(b) and (d) illustrate similar trends are observed on R2R, further emphasizing the impressive generalization power of the partial pretraining strategy.
+5(b) and (d) illustrate similar trends are observed on R2R, further emphasizing the generalization of the partial pretraining strategy.
+ID DSP TSU HRA SR SPL RGSPL
+1 46.98||||52.51 33.73||||36.06 23.03||||22.06
+2 52.00||||52.53 39.21||||39.53 26.54||||24.97
+3 53.17||||53.89 38.31||||40.89 26.73||||26.43
+4 55.07||||51.97 39.50||||37.87 26.70||||25.20
+5 51.75||||54.39 38.86||||41.54 26.56||||27.31
+Table 6: Ablation study on different components of SUSA on the REVERIE (validation||||test) unseen split.
+4) Overall Design.
+To thoroughly evaluate the efficacy of key components, we conduct ablation experiments on the REVERIE validation and test unseen split in the Tab.
+6.
+As shown in Tab.
+6, an intriguing phenomenon is observed: directly incorporating the DSP and TSU modules (#4) results in an improvement on validation unseen (e.g., SR=55.07, SPL=39.50), while performance on test unseen deteriorates.
+Conversely, contrastive learning within the HRA module (#5) produces opposite outcomes, particularly with respect to the SR metric.
+This may be attributed to the learnable token $v$, which, while pooling hybrid environmental representations, also introduces slight noise that causes fluctuations of navigation performance.
+Nevertheless, the overall performance is progressively improved with the inclusion of DSP, TSU and HRA modules.
+Analysis of the HRA module and the computational efficiency of the entire model is presented in supplementary materials.
 4.4 Qualitative Analysis
@@
 This manifest that, thanks to the salient landmarks provided by the SUSA architecture, the agent is better able to navigate to the target location by grounding textual instructions.
-Figure 6: Predicted trajectories of SUSA and DEUT on the R2R validation unseen split.
+Figure 6: Predicted trajectories of SUSA and DUET on the R2R validation unseen split.
 5 Conclusion
-We present the SUSA architecture, which offers complementary environmental representations beyond RGB images.
-We propose the TSU module, enabling the agent to identify the most instruction-relevant views through both static and dynamic matching.
-The DSP module enhances spatial awareness by integrating depth exploration maps.
-We further leverage contrastive learning to precisely align hybrid environmental representations with instructions.
+We present the SUSA architecture, which exploits complementary semantic-spatial representations beyond RGB to facilitate environmental understanding and instruction grounding.
+Specifically, we propose the TSU module, enabling the agent to identify the most instruction-relevant semantic.
+The DSP module enhances spatial awareness by grounding depth exploration maps.
+We further employ contrastive learning to precisely align hybrid environmental representations with instructions.
 Experimental results on the R2R, REVERIE, and SOON benchmarks demonstrate the effectiveness and generalization of our approach.
-Future work will enrich beneficial environmental representations to advance specific VLN tasks.
+Future work will enrich informative environmental representations to advance VLN tasks in continuous environments.
 Supplementary Material
@@
 SOON is another goal-oriented dataset, containing about 5k instructions, each averaging 47 words in length.
-We evaluated the experimental results of the unseen test split on the public online leaderboards***https://eval.ai/web/challenges/challenge-page/606/participate, https://eval.ai/web/challenges/challenge-page/97/participate.
-We assure the readers that we will make our code and model checkpoints publicly available at https://github.com/HCI-LMC/VLN-SUSA.
+We evaluated the experimental results of the unseen test split on the public online leaderboards [^1] https://eval.ai/web/challenges/challenge-page/606/participate, https://eval.ai/web/challenges/challenge-page/97/participate.
+We assure the readers that we will make our code and model checkpoints publicly available.
 7 Additional Experiments
+Validation Unseen
+Methods TL SR SPL RGS RGSPL
+BEVBert [1] - 51.78 36.37 34.71 24.44
+GridMM [56] 23.20 51.37 36.47 34.57 24.56
+SEAT[53] - 49.45 35,51 32.83 23.14
+SUSA(DP) 24.43 52.54 37.59 35.47 25.27
+Depth SUSA(DEM) 22.88 52.00 39.21 35.05 26.54
+KESU [15] 22.93 50.36 35.15 35.02 24.17
+KERM [31] 21.85 50.44 35.38 34.51 24.45
+AACL [33] 22.11 49.42 33.54 33.31 22.49
+CONSOLE [34] - 50.07 34.40 34.05 23.33
+Text SUSA(TP) 23.40 53.17 38.31 37.18 26.73
+Table 9: Comparison with related works on the REVERIE validation unseen split.
 7.1 Comparison with Previous Approaches
-In this subsection, we discuss the differences between the proposed SUSA architecture and previous approaches incorporating depth or textual environmental information.
-For depth-based methods, such as BEVBert [1] and GridMM [58], depth information is primarily leveraged to map the positions of the corresponding RGB images within the environment, facilitating the construction of local or global grid-based representations.
-In contrast, our approach directly incorporates depth images as inputs, providing the agent with a clearer perception of the environmental layout.
-On the other hand, CONSOLE [30] leverages ChatGPT https://openai.com/blog/chatgpt to generate landmark cooccurrence priors, which are then aligned with the agent’s visual observations, while KERM [26] describes the attributes and relationships between objects as textual knowledge.
-Despite the varied forms of incorporating textual semantics, these methods typically compact textual semantics with corresponding visual views before feeding them to the agent.
-In contrast, our approach treats the generated textual semantics as a separate environmental representation, enabling the agent to better ground environments with textual instructions, thereby significantly improving both navigation success and object recognition accuracy.
-In summary, compared to previous approaches, we improve spatial-semantic perception for agent by equipping agent with more alternative environmental representations.
-Figure 7: The weights for hybrid representations in the HRF module change during training (see Eq. 6).
+We demonstrate the effectiveness of the proposed method through comparisons with related approaches.
+As shown in  Tab.
+9, experiments reveal that directly grounding depth information—whether using depth panoramas (DP) or depth exploration maps (DEM) as input—outperforms prior depth-based methods [1, 56, 53].
+Notably, using depth exploration maps leads to better navigation performance than depth panoramas.
+Similarly, directly aligning textual semantics with instructions via textual panoramas (TP) surpasses methods [15, 31, 34] that integrate textual semantics into environmental representations.
+This underscores the superiority of independently grounding each modality, improving the interpretability of each modality’s contribution.
+Figure 7: The weights for hybrid representations in the HRA module change during training (see Eq. 6).
 Figure 8: Similarity matrix for the top 50 most frequent landmarks: REVERIE instructions vs. environments (left), R2R instructions vs. environments (right). Best viewed in color.
@@
 We quantitatively illustrate the weights (see Eq.
-6) assigned to each branch of the HMF module during training (in Fig.
+6) assigned to each branch of the HRA module during training (in Fig.
 7) and inference (in Tab.
-9).
+10).
 These weights are normalized between 0 and 1, with higher values denoting stronger dependency on the corresponding environmental representation.
@@
 6) of hybrid representations during inference on R2R.
-$\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ denotes the fusion weight of textual semantic, local RGB, depth, global RGB representations in the HRF module, respectively.
+$\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ denotes the fusion weight of textual semantic, local RGB, depth, global RGB representations in the HRA module, respectively.
 7.3 Quality of Generated Salient Landmarks
@@
 Notably, the plug-and-play CLIP text encoder excels in understanding image-related descriptions, while BERT is more adept at comprehending natural language sentences.
-Therefore, we use the CLIP text encoder to extract features from the generated salient landmarks, instead of BERT.
-Figure 9: Analysis of the contrastive loss weight $\lambda_{2}$ in HRF module on R2R validation unseen split.
+Therefore, we use the CLIP text encoder to extract features from the generated salient landmarks.
+Figure 9: Analysis of the contrastive loss weight $\lambda_{2}$ in HRA module on R2R validation unseen split.
 7.4 Analysis of the Contrastive Loss Weight
 Our ablation studies indicate that the contrastive learning loss weight, denoted as $\lambda_{2}$ in Eq.
-11, significantly affects model navigation performance.
+11, affects model navigation performance.
 As shown in Fig.
-9, we observe that, within a certain range, increasing the contrastive loss weight improves navigation performance.
+9, agent achieves the optimal navigation performance when the $\lambda_{2}$ is set to 0.8.
+we further observe that, within a certain range, increasing the contrastive loss weight improves navigation performance.
 However, when the contrastive loss weight is too high, performance (SR/OSR) declines, thus overemphasizing contrastive learning loss may lead the agent to prioritize modality alignment over core sequential navigation tasks.
-ID DSP TSU HRF SR SPL RGSPL
-1 46.98/52.51 33.73/36.06 23.03/22.06
-2 52.00/52.53 39.21/39.53 26.54/24.97
-3 53.17/53.89 38.31/40.89 26.73/26.43
-4 55.07/51.97 39.50/37.87 26.70/25.20
-5 51.75/54.39 38.86/41.54 26.56/27.31
-Table 10: Ablation study on different components of SUSA on the REVERIE (validation/test) unseen split.
-7.5 Overall Design on REVERIE
-We further supplemented the performance of each key component of SUSA on the REVERIE task.
-To ensure fairness, we adopt the same hyperparameter setting ($\lambda_{2}$=0.8) as used in R2R, though this may be suboptimal for REVERIE.
+[…]
+1 13.41 3.30 79.86 71.43 60.63
+2 13.85 3.08 80.42 72.80 62.71
+3 13.78 3.06 81.65 72.92 63.16
+4 12.18 3.06 79.99 73.05 64.85
+5 14.58 3.08 82.08 73.95 62.84
+Table 11: Ablation study on different components of SUSA on the R2R validation unseen split.
+7.5 Overall Design on R2R
+We further supplemented the performance of each key component of SUSA on the R2R task.
 As shown in Tab.
-10, our previous conclusions are reaffirmed: the DSP module allows the agent to better perceive spatial layouts, thus improving navigation efficiency (SPL).
-The TSU module enhances the agent’s understanding of textual semantics, resulting in higher success rates in both navigation (SR) and object recognition (RGS).
-An intriguing phenomenon, as shown in Tab.
-10, is observed on the REVERIE dataset: directly incorporating the DSP and TSU modules (#4) results in an improvement on validation unseen (e.g., SR=55.07, SPL=39.50), while performance on test unseen deteriorates.
-In contrast, the addition of the HRF module (#5) yields opposite results.
-Overall, the introduction of the HRF module leads to superior navigation performance, both in terms of navigation (SR/SPL) and object grounding (RGSPL).
+6, our previous conclusions are reaffirmed: the DSP module allows the agent to better perceive spatial layouts, thus improving navigation efficiency (SPL).
+The TSU module is better at understanding textual semantics, resulting in higher success rates.
+After integrating the DSP and TSU modules, the agent’s performance improved significantly (e.g., SR=73.05, SPL=64.85).
+When the HRA module is further introduced, the navigation success rate on R2R increases to 73.95, indicating that contrastive learning further enhances instruction-environment alignment.
+However, its SPL metric remains suboptimal, suggesting that the contrastive learning strategy in the HRA module may excel at accurately aligning the environment with instructions but lacks the capability to optimize the navigation path length.
 7.6 Computing Efficiency vs. Navigation Performance
@@
 Compared to the baseline DUET, our approach introduces spatial and semantic environmental information, which inevitably increasing computational cost.
-However, despite this modest cost, the SUSA architecture notably enhances agent navigation performance.
+However, benefiting from the partial pretraining strategy and a shared panorama encoder for both the depth and RGB environmental information, our SUSA architecture notably enhances agent navigation performance with modest computational resources.
+The frozen BLIP-2 and CLIP models provide the agent with off-the-shelf textual semantics and are excluded from the training or inference.
 8 Qualitative Examples
@@
 2) Beyond landmarks, we will construct direction-related textual semantics (e.g., “past,” “turn left”) to better model the spatial relationships between actions in instructions and the environmental layout or historical exploration maps.
-We also intend to investigate how well our approach generalizes to other navigation tasks.
+In the future, we also intend to explore the generalizability of our approach, which integrates richer representations from multiple perspectives, in embodied tasks such as continuous navigation and embodied question answering.
 Additionally, given the complexity of VLN tasks, current mainstream VLN frameworks are often intricate and cumbersome, with a wide variety of evaluation metrics that are difficult to balance.

### v5

@@
-Agent Journey Beyond RGB: Unveiling Hybrid Semantic-Spatial Environmental Representations for Vision-and-Language Navigation
+Agent Journey Beyond RGB: Hierarchical Semantic-Spatial Representation Enrichment for Vision-and-Language Navigation
 Abstract
 Navigating unseen environments based on natural language instructions remains difficult for egocentric agents in Vision-and-Language Navigation (VLN).
-Existing approaches primarily rely on RGB images for environmental representation, underutilizing latent textual semantic and spatial cues and leaving the modality gap between instructions and scarce environmental representations unresolved.
-Intuitively, humans inherently ground semantic knowledge within spatial layouts during indoor navigation.
-Inspired by this, we propose a versatile Semantic Understanding and Spatial Awareness (SUSA) architecture to encourage agents to ground environment from diverse perspectives.
-SUSA includes a Textual Semantic Understanding (TSU) module, which narrows the modality gap between instructions and environments by generating and associating the descriptions of environmental landmarks in agent’s immediate surroundings.
-Additionally, a Depth-enhanced Spatial Perception (DSP) module incrementally constructs a depth exploration map, enabling a more nuanced comprehension of environmental layouts.
-Experiments demonstrate that SUSA’s hybrid semantic-spatial representations effectively enhance navigation performance, setting new state-of-the-art performance across three VLN benchmarks (REVERIE, R2R, and SOON).
-The source code will be publicly available.
+Intuitively, humans inherently ground concrete semantic knowledge within spatial layouts during indoor navigation.
+Although previous studies have introduced diverse environmental representations to enhance reasoning, other co-occurrence modalities are often naively concatenated with RGB features, resulting in suboptimal utilization of each modality’s distinct contribution.
+Inspired by this, we propose a hierarchical Semantic Understanding and Spatial Awareness (SUSA) architecture to enable agents to perceive and ground environments at diverse scales.
+Specifically, the Textual Semantic Understanding (TSU) module supports local action prediction by generating view-level descriptions, thereby capturing fine-grained environmental semantics and narrowing the modality gap between instructions and environments.
+Complementarily, the Depth-enhanced Spatial Perception (DSP) module incrementally constructs a trajectory-level depth exploration map, providing the agent with a coarse-grained comprehension of the global spatial layout.
+Extensive experiments demonstrate that SUSA’s hierarchical representation enrichment not only boosts the navigation performance of the baseline on discrete VLN benchmarks (REVERIE, R2R, and SOON), but also exhibits superior generalization to the continuous R2R-CE.
+Codehttps://github.com/HCI-LMC/VLN-SUSA
 1 Introduction
-With the advancement of multimodal technology [44, 47, 29, 59], several embodied VLN tasks [2, 41, 37, 8] have garnered significant attention due to their promising applications in robotics and intelligent assistance.
-Among them, conventional VLN tasks [2, 27] focused on step-by-step navigation within unseen environments based on detailed instructions, while goal-oriented VLN tasks [64, 41, 45] demand agents to identify predefined objects based on high-level instructions.
-Consequently, VLN agents are required to profoundly cognize the spatial information of environment and ground them with textual instructions to brilliantly accomplish arduous navigation tasks.
+With the advancement of multimodal technology (Radford et al.
+2021; Li et al.
+2023a; Zhang et al.
+2024) in the artificial intelligence community, embodied VLN tasks (Krantz et al.
+2020; Qi et al.
+2020; Das et al.
+2018) have garnered significant attention due to their promising applications in robotics and intelligent assistance.
+Conventional VLN tasks (Anderson et al.
+2018; Krantz et al.
+2020) focused on step-by-step navigation in unseen environments based on detailed instructions, while goal-oriented VLN tasks (Zhu et al.
+2021; Qi et al.
+2020) demand agents to identify predefined objects based on high-level instructions.
+Consequently, VLN agents are required to profoundly cognize the environmental spatial information and ground them with textual instructions to brilliantly accomplish navigation tasks.
 Figure 1: An overview of the proposed SUSA.
-Unlike previous approaches that solely rely on RGB input, we introduce view-level textual panoramas and trajectory-level depth exploration maps, supporting the agent’s explicit understanding of the environment.
-Considerable literature has exhibited the efficacy of Transformer-based models [5, 20, 38, 52, 40] for interacting with instructions, RGB-based environmental observations, and historical trajectories.
-However, we suggest that RGB-only visual environment representations leave VLN fraught with two challenges: the modality gap between instructions and the environment, and the scarcity of training environmental representations.
-Firstly, VLN agents struggle to precisely match the landmarks in language instructions with the entities in visual environments, due to the intrinsic modality heterogeneity.
-To address this, recent works [12, 32] utilized contrastive learning to promote alignment, while others leveraged candidate object concepts [50, 33], large language models (LLMs) [34] or retrieved knowledge [31] to explicitly reason the textual semantics (i.e., linguistic descriptions about landmarks) to enhance environmental representations.
-However, these methods do not directly align textual semantics with landmarks (e.g., “fireplace” in Fig.
-1) in instruction, yielding trivial gain.
-The second challenge pertains to the monotonous and limited environmental representations, which often leads agents to overfit to familiar spatial layouts [51].
-To alleviate this limitation, recent endeavors [55, 7] augment the quantity of environmental scenes, while others diversify environmental representations by constructing graphs [16, 23], grids [56, 1], or volumetric environments [36].
-However, these RGB-based methods easily suffer from attention imbalance or environmental bias [65, 61].
-FDA [18] utilizes frequency information to enrich environmental representations.
-However, depth images vividly convey spatial information (e.g., doors and corridors), which is often overlooked in discrete navigation tasks [49].
-Few works [1, 56] incorporate depth, they merely project depth values onto the corresponding spatial locations of RGB images, rather than independently grounding depth[53].
-Overall, the underlying textual semantic and spatial representations remains underutilized, though studies [22] suggest that injecting inductive biases where different modules access distinct modalities, offers a promising solution.
-In this work, we propose the Semantic Understanding and Spatial Awareness (SUSA) architecture, which individually grounds semantic and spatial representations of the environment with instructions and synergistically predicts navigation actions from diverse perspectives.
-As illustrated in Fig.
-1, we first construct a view-level textual panorama by extracting salient landmarks as the textual semantics to explicitly match them with language instructions, and select the most instruction-relevant view, thereby bridging the semantic gap between the visual and linguistic modalities.
-Further, beyond RGB images, we progressively construct an independent trajectory-level depth exploration map to enhance the agent’s spatial perception of the environment, also directly aligning it with instructions.
-To enable the agent to absorb complementary information from different environmental perspectives, we aggregate these hybrid semantic-spatial environmental representations to synergistically facilitate hierarchical navigation predictions, and employ contrastive learning to align these hybrid representations with language instructions.
-We empirically substantiate that exclude semantic panoramas and depth exploration maps during pretraining manifest markedly enhanced generalization capabilities.
-Experimental results demonstrate that our agent achieves significantly improved navigation performance, outperforming state-of-the-art methods across three vision-and-language navigation benchmarks: R2R [2], REVERIE [41], and SOON [64].
-The main contributions of our SUSA are as follows:
-  • SUSA devises a textual-aware semantic understanding (TSU) module to match the most related landmarks between environments and instructions, explicitly addressing the modality heterogeneity.
-  • SUSA further presents a depth-enhanced spatial perception (DSP) module to explore and represent environmental layouts beyond surface RGB content.
-  • Extensive experiments on three VLN benchmarks have demonstrated the superiority of hybrid semantic-spatial environmental representations introduced by SUSA.
+Beyond RGB inputs, we introduce view-level textual panoramas and trajectory-level depth exploration maps, supporting the agent’s explicit understanding of the environment.
+Recently, increasing the diversity and quantity of augmented environments (Wang et al.
+2023b, c; Liu, Wang, and Yang 2024) has emerged as a popular strategy to facilitate the generalization ability of VLN agents in unseen scenarios.
+This trend is largely motivated by the limited navigational environments in simulators, as well as the powerful fitting capacity of Transformer-based VLN models (Chen et al.
+2021; Pashevich, Schmid, and Sun 2021).
+Nonetheless, VLN agents still struggle to accurately ground the landmarks mentioned in language instructions with the RGB-based visual scene, due to the intrinsic modality heterogeneity.
+To address this, recent studies introduce textual semantics as perceptual representations to guide navigation (Pan et al.
+2024).
+These approaches leverage candidate object concepts (Wang et al.
+2023a; Lin et al.
+2023), large language models (LLMs) (Lin et al.
+2024), or external knowledge (Li et al.
+2023b) to explicitly reason about textual descriptions of landmarks and enhance environmental grounding.
+However, these methods tend to combine textual and visual representations in a straightforward manner, which limits the effective utilization of each modality and constrains the interpretability of their individual contributions.
+Similarly, co-occurring environmental cues from other modalities—such as frequency (He et al.
+2024), spectrum (Hwang et al.
+2023), and depth (Tan et al.
+2022)—have also been preliminarily explored, though remain underutilized.
+In this work, we propose the hierarchical Semantic Understanding and Spatial Awareness (SUSA) architecture, which augments the agent’s environmental comprehension by effectively perceiving and exploiting fine-grained textual semantics and coarse-grained depth-aware spatial information.
+To enable fine-grained semantic grounding, as shown in Fig.
+1, we construct a view-level textual panorama by extracting salient landmarks as textual semantics, thereby bridging the gap between visual and linguistic modalities.
+The most instruction-relevant view is then selected for local navigation prediction.
+For global action prediction, we progressively construct a trajectory-level depth exploration map to strengthen the agent’s coarse-grained spatial perception.
+Finally, SUSA hierarchically aggregates these hybrid semantic-spatial representations to perceive complementary environmental information and employs contrastive learning to align them with natural language instructions.
+Extensive experiments on four VLN benchmarks (R2R (Anderson et al.
+2018), REVERIE (Qi et al.
+2020), SOON (Zhu et al.
+2021) and R2R-CE (Krantz et al.
+2020)) demonstrated the superiority of hybrid semantic-spatial representations enrichment.
+The contributions of our SUSA are as follows:
+  • SUSA devises a textual-aware semantic understanding (TSU) module to perceive fine-grained textual semantics of environments, and explicitly address modality heterogeneity between instructions and environments.
+  • SUSA further constructs a depth exploration map within the depth-enhanced spatial perception (DSP) module to complementarily perceive global environmental layouts and reinforce holistic spatial awareness beyond RGB.
+  • Experiments demonstrate that SUSA substantially improves upon the baseline, achieving state-of-the-art performance on three discrete VLN tasks, and exhibits promising generalization in continuous environments.
 2 Related Work
 Vision-and-Language Navigation (VLN).
-Vision-and-Language Navigation (VLN) tasks [2, 41, 64] require an agent to navigate in embodied environments sequentially towards a designated location or object based on natural language instructions.
-Given that the VLN task can be formulated as a partially observable Markov decision process, early methods [54, 66, 14, 26] primarily relied on reinforcement learning and imitation learning paradigms to improve navigation performance, incorporating various strategies and auxiliary loss [63, 54] designs.
-More recently, to better generalize to unknown environments and ground instructions, many researchers [5, 20, 38, 52, 40] have focused on transformer-based architectures, which effectively model interactions among instructions, environmental features, and historical navigation trajectories.
-While these approaches have propelled VLN forward, they mainly rely on fragile RGB-based environment representations.
-This paper encourages the agent to perceive and ground the environment from text semantics and depth perspectives.
-Semantic Alignment in VLN.
-To align the semantic relationship between the environment and instructions, current practitioners in the VLN community can be divided into two categories: implicit and explicit approaches.
-Implicit methods [12, 51, 32] commonly employ contrastive learning or attention mechanisms to associate visual contexts with textual instructions.
-However, given the inherent semantic gap between these modalities, implicit approaches have limited alignment capabilities.
-Explicit alignment techniques aim to project both semantics into a same modality space.
-For example, some methods transform the environment into a text-based modality, Console [34] scores landmark co-occurrences to identify key objects, and AACL [33] leverages CLIP [44] to refine textual entities.
-Similar to our approach, KERM [31] utilizes retrieved knowledge as textual semantics for each view.
-However, these explicit methods directly “stitch” textual and visual environment representations, which may lead to an overreliance on text modality yet underutilizing visual data [67].
-We thus take environmental textual semantic as an independent modality to directly match the most instruction-relevant landmarks.
+Vision-and-Language Navigation (VLN) tasks (Anderson et al.
+2018; Qi et al.
+2020; Zhu et al.
+2021) require an agent to navigate in embodied environments sequentially towards a designated location or object based on natural language instructions.
+Given that the VLN task can be formulated as a partially observable Markov decision process, early methods (Wang et al.
+2019; Fried et al.
+2018) primarily relied on reinforcement learning and imitation learning paradigms to improve navigation performance, incorporating various strategies and auxiliary loss (Wang et al.
+2019) designs.
+More recently, to better model the relationship between the environment and instructions, many researchers (Chen et al.
+2021; Pashevich, Schmid, and Sun 2021) have focused on transformer-based architectures, which effectively model interactions among instructions, environmental features, and historical navigation trajectories.
 Environment Representations in VLN.
 Limited training sources and environmental representations in VLN tasks hinder the agent to generalize the unknown layouts.
-Prior endeavors mitigate this limitation through data augmentation [35, 55] or by diversifying visual representations [60, 1, 56, 36].
-However, these approaches primarily rely on RGB images, while alternative modalities such as frequency [18], spectral[24], depth [49] and semantic information [53, 21, 4] remain relatively underexplored.
-Among them, SEAT [53] enhances cross-domain environmental perception by querying the Matterport3D simulator[3] for depth and semantics.
+To alleviate this limitation, recent endeavors (Wang et al.
+2023b; Chen et al.
+2022a) augment the quantity of environmental scenes, while others diversify environmental representations by constructing graphs (Georgakis et al.
+2022; Huang et al.
+2023), grids (Wang et al.
+2023c; An et al.
+2023), or volumetric environments (Liu, Wang, and Yang 2024).
+Beyond RGB inputs, recent efforts have explored other co-occurrence multimodal(Li et al.
+2023b; He et al.
+2024) cues to enrich environmental representations.
+For instance, SEAT (Wang et al.
+2024b) enhances cross-domain environmental perception by querying the Matterport3D simulator (Chang et al.
+2017) for depth and semantics.
 However, SEAT fuses these multi-type representations into a unified space without explicitly grounding each modality.
-Moreover, these semantic representations [53, 21] are not inherently textual, limiting their ability to bridge the modality gap.
-In contrast, our method constructs the view-level textual panoramas and trajectory-level depth exploration maps, directly grounding them with instructions to perceive rich environmental representations.
+Moreover, these semantic representations (Wang et al.
+2024b; Hong et al.
+2023) are not inherently textual, limiting their ability to bridge the modality gap.
+Other approaches incorporate depth information (An et al.
+2023; Wang et al.
+2023c); however, they typically project depth values onto the spatial coordinates of RGB images without independently grounding the depth modality.
+In contrast, our method constructs the view-level textual panoramas and trajectory-level depth exploration maps, individually grounding them with instructions to perceive rich environmental representations.
 Figure 2: The detailed architecture of our proposed SUSA model.
 The orange and green arrows highlight our proposed Textual-Aware Semantic Understanding (TSU) and Depth-Aware Spatial Perception (DSP) modules, respectively.
-The Hybrid Representation Aggregation (HRA) module is designed for aligning environmental representations with instructions.
+The Hierarchical Aggregation and Prediction (HAP) module is designed for hierarchical aligning environmental representations with instructions.
 3 Methodology
@@
 Task Formulation.
-We focus on discrete VLN tasks [2, 41, 64] within a Matterport3D simulator [3], modeled as an undirected graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}$ denotes nodes and $\mathcal{E}$ signifies the connectivity paths.
+We focus on discrete VLN tasks (Anderson et al.
+2018; Qi et al.
+2020; Zhu et al.
+2021) within a Matterport3D simulator (Chang et al.
+2017), modeled as an undirected graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}$ denotes nodes and $\mathcal{E}$ signifies the connectivity paths.
 At the outset of each episode, the agent equipped with a GPS sensor, RGB and depth cameras, is situated at a randomly navigable node and receives a language instruction $I=\{\text{w}_{i}\}_{i=1}^{l}$ including $l$ words.
-We use BERT [10] to extract instruction features and then embed them as =$\mathcal{I}=\{w_{1},\dotso,w_{l}\}$.
+We use BERT (Devlin et al.
+2019) to extract instruction features and then embed them as $\mathcal{I}=\{w_{1},\dotso,w_{l}\}$.
 At each time step $t$, the agent can perceive the surrounding RGB panorama $R_{t}=\{\text{r}_{t,i}\}_{i=1}^{n=36}$ and depth panorama $D_{t}=\{\text{d}_{t,i}\}_{i=1}^{n=36}$ along with $n$ corresponding views.
-During navigation, the agent make the next action $a_{t}$ by selecting one navigable node from the candidate actions $\mathcal{A}_{t}=\{a_{t,i}\}_{i=0}^{k}$.
-The sequential decision-making process is iterated until the agent decides to stop at the current location or exceeds the maximum navigation steps.
-For goal-oriented tasks [41, 64], the agent is also tasked with locating a target object.
+During navigation, the agent makes the next action $a_{t}$ by selecting one navigable node from the candidate actions $\mathcal{A}_{t}=\{a_{t,i}\}_{i=0}^{k}$.
 Overview of the Proposed Method.
-Our objective is to accord complementary environmental content by equipping the agent with a textual semantic panorama and a depth exploration map.
-Following prevailing approaches [18, 30, 19], we adopt DUET [6] as our baseline, which merely utilizes RGB panoramas and topological maps as environmental representations for both local and global action predictions.
-The SUSA framework consists of three main modules: (a) textual-aware semantic selection is detailed in Sec.
-3.2, (b) depth-aware visual environment exploration is presented in Sec.
-3.3, and (c) hybrid representation aggregation and hierarchical action prediction is subsequently introduced in Sec.
-3.4, as illustrated in Fig.
+Our objective is to enrich complementary environmental content by equipping the agent with a textual semantic panorama and a depth exploration map.
+Following prevailing approaches (He et al.
+2024), we adopt DUET (Chen et al.
+2022b) as our baseline, which merely utilizes RGB panoramas and topological maps as environmental representations for both local and global action predictions.
+The SUSA framework consists of three main modules: (a) TSU: textual-aware semantic selection is detailed in Section 3.2, (b) DSP: depth-enhanced visual environment exploration is presented in Section 3.3, and (c) HAP: hybrid representation aggregation and hierarchical action prediction is subsequently introduced in Section 3.4, as illustrated in Fig.
 2.
@@
 In this module, we build a textual semantic panorama that enables the agent to understand the environment from a textual perspective, allowing it to select the most relevant navigable view as the optimal local navigation action.
-Figure 3: Illustration of TSU module in  Fig. 2, which matches instructions and semantic features by static and dynamic strategies.
+Figure 3: Illustration of the TSU module in Fig.2, which matches instructions and semantic features by static and dynamic matching strategies.
 Salient Landmarks Extractor.
@@
 As illustrated in Fig.
-2, we utilize off-the-shelf BLIP-2-FlanT5-xxL [29], a powerful visual-language model, to discretize $R_{t}$ and generate corresponding textual descriptions for each view using the simple prompt “a photo of”.
-Following [15, 28], this process yields a textual panorama $\mathcal{S}_{t}=\{s_{t,i}\}_{i=1}^{n=36}$, where each view $s_{t,i}$ contains informed descriptions of salient landmarks.
+2, we utilize off-the-shelf BLIP-2-FlanT5-xxL (Li et al.
+2023a), a powerful visual-language model, to discretize $R_{t}$ and generate corresponding textual descriptions for each view using the simple prompt “a photo of”.
+Following (Gao et al.
+2024; Li and Bansal 2023), this process yields a textual panorama $\mathcal{S}_{t}=\{s_{t,i}\}_{i=1}^{n=36}$, where each view $s_{t,i}$ contains informed descriptions of salient landmarks.
 These descriptions typically involve room types and objects (e.g.,“a living room with two chairs”), which assist in the object recognition.
 We further discuss the quality of the generated descriptions in supplementary materials.
-Additionally, we employ the CLIP (ViT-L-14-336px) text encoder [44] to extract textual semantic panorama features, embedding them as $\mathcal{X}^{\mathcal{S}}_{t}=\{x^{s}_{t,1},\dots,x^{s}_{t,n}\}$.
+Additionally, we employ the CLIP (ViT-L-14-336px) text encoder (Radford et al.
+2021) to extract textual semantic panorama features, embedding them as $\mathcal{X}^{\mathcal{S}}_{t}=\{x^{s}_{t,1},\dots,x^{s}_{t,n}\}$.
 Thanks to the powerful yet frozen BLIP-2 and CLIP models, we empower the agent to obtain textual semantic features from the navigable nodes in its vicinity during navigation.
 Textual Semantic Selector.
-Previous methods [50, 33, 31] aligned environmental RGB images with instructions by exploiting textual semantics yet the inherent modality discrepancy between visual and language modalities remained unaddressed.
-Similar to [39], we treat the textual semantic as a perceptual representation for navigation in this module, directly matching salient landmarks with the instructions.
+Similar to (Pan et al.
+2024), we treat the textual semantic as a perceptual representation for navigation in this module, directly matching salient landmarks with the instructions.
 As shown in Fig.
 2 and Fig.
-3, our textual semantic selector combines both $static$ and $dynamic$ matching strategies to select the most instruction-relevant view in the textual panorama $\mathcal{X}^{\mathcal{S}}_{t}$.
+3, our textual semantic selector combines both staticstatic and dynamicdynamic matching strategies to select the most instruction-relevant view in the textual panorama $\mathcal{X}^{\mathcal{S}}_{t}$.
 Static Matching: The most straightforward approach to semantic alignment is by calculating similarity, where higher similarity indicates stronger relevance between the textual semantics of each view and the instruction.
@@
 Therefore, we then apply row-wise max-pooling to purify the static instruction relevance $\Gamma_{t}^{stat}\in\mathbb{R}^{n}$ to match the most relevant landmarks between textual semantics and instructions at step $t$ :
-$\small\begin{split}\mathbf{M}_{i,j}=\frac{\mathcal{X}^{\mathcal{S}}_{t,i}\cdot\mathcal{I}_{j}}{\|\mathcal{X}^{\mathcal{S}}_{t,i}\|\|\mathcal{I}_{j}\|},\ \ \ \ \Gamma_{t}^{stat}=\max_{j}\mathbf{M}_{i,j}\end{split}$ (1)
+$\begin{split}\mathbf{M}_{i,j}=\frac{\mathcal{X}^{\mathcal{S}}_{t,i}\cdot\mathcal{I}_{j}}{\|\mathcal{X}^{\mathcal{S}}_{t,i}\|\|\mathcal{I}_{j}\|},\ \ \ \ \Gamma_{t}^{stat}=\max_{j}\mathbf{M}_{i,j}\end{split}$ (1)
 where $i=1,\dots,n$ and $j=1,\dots,l$.
@@
 Specifically, we utilize a standard multi-layer transformer decoder to dynamically match the relevance $\Gamma_{t}^{dyn}$ between textual semantics and instructions, as formalized by:
-$\small\begin{split}\Gamma_{t}^{dyn}=\text{TCA}(\mathcal{X}^{\mathcal{S}}_{t},\mathcal{I},\mathcal{I}),\end{split}$ (2)
+$\begin{split}\Gamma_{t}^{dyn}=\text{TCA}(\mathcal{X}^{\mathcal{S}}_{t},\mathcal{I},\mathcal{I}),\end{split}$ (2)
 where TCA denotes a textual modality cross-attention mechanism that models textual semantic-instruction relationships.
@@
 3.3 Depth-Enhanced Spatial Perception
-Depth images, which also contain more intuitive and easily distinguishable structural information, were not adequately exploited in earlier works.
-To enhance the spatial perception, we concurrently construct a depth-based exploration map to memorize depth-based navigation trajectory alongside the RGB-based one [6].
+Depth images, which also contain more intuitive and easily distinguishable structural information, were not adequately exploited in earlier works (Tan et al.
+2022).
+To enhance the spatial perception, we concurrently construct a depth-based exploration map to memorize depth-based navigation trajectory alongside the RGB-based one (Chen et al.
+2022b).
 Depth-Enhanced Exploration Mapping.
@@
 As illustrated in Fig.
-2, we then employ ResNet-50 [17] (pretrained on the Gibson [57]) and CLIP (ViT-L/14@336px) as the depth and RGB encoders, respectively, to extract panoramic depth features $\mathcal{D}_{t}\in\mathbb{R}^{n\times d_{v}}$ and panoramic RGB features $\mathcal{R}_{t}\in\mathbb{R}^{n\times d_{v}}$, where $d_{v}$ represents the dimension of each view.
+2, we then employ ResNet-50 (He et al.
+2016) (pretrained on the Gibson (Xia et al.
+2018)) and CLIP (ViT-L/14@336px) as the depth and RGB encoders, respectively, to extract panoramic depth features $\mathcal{D}_{t}\in\mathbb{R}^{n\times d_{v}}$ and panoramic RGB features $\mathcal{R}_{t}\in\mathbb{R}^{n\times d_{v}}$, where $d_{v}$ represents the dimension of each view.
 Moreover, we employ a two-layer transformer-based panorama encoding module to perform self-attention and sequentially encode the depth and RGB panoramic images.
@@
 Subsequently, we aggregate each visited node and its adjacent views by average pooling to construct depth and RGB exploration maps in parallel.
-The explored depth and RGB trajectory maps are formalized as $\mathcal{T}_{t}^{\mathcal{D}}\in\mathbb{R}^{k\times d_{n}}$ and $\mathcal{T}^{\mathcal{R}}_{t}\in\mathbb{R}^{k\times d_{n}}_{t}$ Here, $k$ denotes the number of navigable nodes at step $t$, while $d_{n}$ represents the dimension of each navigable node.
-These exploration maps memorize the agent navigation trajectory information for global action predictions, which serves to guide the agent traversal across various navigable nodes when considering backtracking.
-Cross-Modal Interaction and Reasoning.
-To defer to instructions, we utilize a multi-layer cross-encoder that interacts with both the language instructions and the depth exploration map features to generate the refined depth map $\tilde{\mathcal{T}^{\mathcal{D}}}$.
-This encoder comprises four residual connection layers of a cross-modal transformer, each incorporating graph-aware self-attention (GASA) [6], cross-modal attention, and a feed-forward network.
-The GASA operation is formally expressed as $\text{GASA}(X)=\eth(X\Theta_{q}(X\Theta_{k})^{T}+\mathcal{M})X\Theta_{v}$ where $\mathcal{M}$ represents a learnable distance matrix encoding the relative distances between navigation nodes, $\eth$ denotes the Softmax activation function, and $\Theta_{q},\Theta_{k},\Theta_{v}$ are learnable weight matrices.
-In parallel, we facilitate similar cross-modal interactions between the language instructions and the RGB exploration map $\mathcal{T}^{\mathcal{R}}$, as well as the RGB panoramas $\mathcal{X}^{\mathcal{R}}$.
-These are encoded as $\tilde{\mathcal{T}^{\mathcal{R}}}$ and $\tilde{\mathcal{X}^{\mathcal{R}}}$, thereby enabling the model to perform both global and local predictions based on the given instructions $\mathcal{I}$.
-Compared to the standalone $\mathcal{T}^{\mathcal{R}}$, the incorporation of a depth exploration map $\mathcal{T}^{\mathcal{D}}$ enhances the perceptual clarity of environmental structures, mitigating the risk of overfitting to RGB-based visual noises [25] or biases [22, 51].
-Figure 4: Illustration of the pipeline for fusing hybrid environmental representations in the proposed HRA module in Fig. 2.
-3.4 Hybrid Representation Aggregation
+The explored depth and RGB trajectory maps are formalized as $\mathcal{T}_{t}^{\mathcal{D}}\in\mathbb{R}^{k\times d_{n}}$ and $\mathcal{T}^{\mathcal{R}}_{t}\in\mathbb{R}^{k\times d_{n}}_{t}$ .
+Here, $k$ denotes the number of navigable nodes at step $t$, while $d_{n}$ represents the dimension of each navigable node.
+These exploration maps memorize the agent navigation trajectory information for global action predictions, which serve to guide the agent traversal across various navigable nodes when considering backtracking.
+[…]
+Figure 4: The pipeline for fusing hybrid environmental representations in the proposed HAP module in Fig. 2.
+3.4 Hierarchical Aggregation and Prediction
 Building on aforementioned methods, at the current state $t$, the agent synchronously perceives the environment through multiple modalities—textual semantic, depth, and RGB —at various scales (local and global).
 To encourage the agent to aggregate complementary context from these hybrid representations, we encode these features into a low-dimensional latent space via attention pooling and apply contrastive learning to align them with the instructions.
-Fusing Hybrid Semantic-Spatial Environmental Representations.
+Hybrid Representations Aggregation.
 As depicted in Fig.
-4, we employ attention pooling to project all environmental representations at time step $t$, along with the corresponding instruction, into an equivalent dimension $d$.
+4, we employ attention pooling to project all environmental representations at step $t$, along with the corresponding instruction, into an equivalent dimension $d$.
 Specifically, the attention pooling AP operation is defined as follows:
-^$\hat{\mathcal{H}}=\text{AP}(\mathcal{H},v),\quad\mathcal{H}\in\{\tilde{\mathcal{X}^{\mathcal{S}}},\tilde{\mathcal{T}^{\mathcal{D}}},\tilde{\mathcal{X}^{R}},\tilde{\mathcal{T}^{R}},\mathcal{I}\},$ (4)
-where $v\in\mathbb{R}^{d}$ denoting the learnable query vector and ^$\hat{\mathcal{H}}\in\mathbb{R}^{d}$. The AP operator is given by:
+$\hat{\mathcal{H}}=\text{AP}(\mathcal{H},v),\quad\mathcal{H}\in\{\tilde{\mathcal{X}^{\mathcal{S}}},\tilde{\mathcal{T}^{\mathcal{D}}},\tilde{\mathcal{X}^{R}},\tilde{\mathcal{T}^{R}}$$\mathcal{I}\},$ (4)
+where $v\in\mathbb{R}^{d}$ denoting the learnable query vector and $\hat{\mathcal{H}}\in\mathbb{R}^{d}$. The AP operator is given by:
 $\eta=\eth\left(\tau(\mathcal{H})\otimes v^{\top}\right),\quad\hat{\mathcal{H}}=\tau\left(\eta\odot\mathcal{H}\right),$ (5)
 with $\tau$ representing the tanh activation function. Subsequently, we perform weighted aggregation of the hybrid environmental representations $E^{\mathrm{hyb}}$, as follows:
-$\small\footnotesize\mathcal{B}=\begin{bmatrix}\beta_{1}&\beta_{2}&\beta_{3}&\beta_{4}\end{bmatrix}^{\top}=\sigma(\mathrm{FFN}[\tilde{\mathcal{X}_{0}^{\mathcal{S}}};\tilde{\mathcal{X}_{0}^{\mathcal{R}}};\tilde{\mathcal{T}_{0}^{\mathcal{D}}};\tilde{\mathcal{T}_{0}^{\mathcal{R}}}])$ (6)
-$\small\footnotesize E^{\mathrm{hyb}}=\begin{bmatrix}\hat{\mathcal{X}^{\mathcal{S}}}&\hat{\mathcal{X}^{\mathcal{R}}}&\hat{\mathcal{T}^{\mathcal{D}}}&\hat{\mathcal{T}^{\mathcal{R}}}\end{bmatrix}\cdot\mathcal{B},$ (7)
-where $\sigma$ stands for the sigmoid activation function, and [;][;][ ; ] denotes concatenation.
+[…]
+$E^{\mathrm{hyb}}=\begin{bmatrix}\hat{\mathcal{X}^{\mathcal{S}}}&\hat{\mathcal{X}^{\mathcal{R}}}&\hat{\mathcal{T}^{\mathcal{D}}}&\hat{\mathcal{T}^{\mathcal{R}}}\end{bmatrix}\cdot\mathcal{B}$, (7)
+where $\sigma$ stands for the sigmoid activation function, and $[;]$ denotes concatenation.
 […]
@@
 $P_{\mathcal{T}}=\beta_{3}P_{\mathcal{T}^{\mathcal{D}}}+\beta_{4}P_{\mathcal{T}^{\mathcal{R}}}.$ (9)
-We convert local predictions into the global action space and predict action $a_{t}$ via dynamic fusing [6] $P_{\mathcal{T}}$ and $P_{\mathcal{X}}$.
-3.5 Training and Inference
-Pretraining.
-Our pre-training phase followed prior work [6, 30], utilizing auxiliary tasks such as Masked Language Modeling (MLM), Masked Region Classification (MRC), and Single-step Action Prediction (SAP).
+We convert local predictions into global actions and predict $a_{t}$ via dynamic fusing (Chen et al. 2022b) $P_{\mathcal{T}}$ and $P_{\mathcal{X}}$.
+3.5 Training
+Partially Pretraining.
+Our pre-training phase followed prior work (Chen et al.
+2022b), utilizing auxiliary tasks such as Masked Language Modeling (MLM), Masked Region Classification (MRC), and Single-step Action Prediction (SAP).
 Object Grounding (OG) task to localize target objects in REVERIE and SOON.
-Given the sparsity of depth images and textual semantics compared to RGB images, completely pre-train all parameter of SUSA may lead to overfitting and increased training costs.
+Given the sparsity of depth images and textual semantics compared to RGB images, completely pre-training all parameter of SUSA may lead to overfitting and increased training costs.
 As shown in Fig.
 2, only the computation flows indicated by black arrows were pre-trained.
-The effectiveness of this partial pre-trained strategy is demonstrated in Sec.
-4.3.
-Methods Validation Seen Validation Unseen Test Unseen
-[…]
-RCM [54] 10.70 23.33 21.82 16.23 15.36 11.98 9.29 6.97 4.89 3.89 10.60 7.84 6.67 3.67 3.14
-HOP+ [43] 10.59 55.87 49.55 40.76 36.22 14.57 36.07 31.13 22.49 19.33 15.17 33.82 28.24 20.20 16.86
-DSRG [50] - 75.69 68.09 61.07 54.72 - 47.83 34.02 32.69 23.37 - 54.04 37.09 32.49 22.18
-GridMM [56] - - - - - 23.20 51.37 36.47 34.57 24.56 19.97 53.13 36.60 34.87 23.45
-AZHP [58] 13.60 70.98 62.24 56.99 50.14 22.08 49.02 36.25 32.41 24.13 21.10 52.52 36.11 32.10 22.54
-FDA [18] - - - - - 19.04 47.57 35.90 32.06 24.31 17.30 49.62 36.45 30.34 22.08
-CONSOLE [34] - 74.14 65.15 60.08 52.69 - 50.07 34.40 34.05 23.33 - 55.13 37.13 33.18 22.25
-KERM [31] 14.25 71.89 64.04 57.55 51.22 21.85 49.02 34.83 33.97 24.14 18.38 52.26 37.46 32.69 23.15
-VER [36] 16.13 75.83 66.19 61.71 56.20 23.03 55.98 39.66 33.71 23.70 24.74 56.82 38.76 33.88 23.19
-DUET [6] 13.86 71.75 63.94 57.41 51.14 22.11 46.98 33.73 32.15 23.03 21.30 52.51 36.06 31.88 22.06
-SUSA (Ours) 14.60 76.95 69.07 61.77 55.86 22.59 51.75 38.86 35.02 26.56 17.86 54.39 41.54 36.11 27.31
-Table 1: Comparison with the state of the art on REVERIE. Bold highlight the best performance in each column, $\uparrow$ indicates better performance with higher values.‘–’: unavailable statistics.
-Fine-tuning and Inference.
-During fine-tuning, we employ a contrastive learning loss, $\mathcal{L}_{cl}$ to align the hybrid environmental representations $E^{hyb}$ with the corresponding instructions $\mathcal{I}$:
+The effectiveness of this partial pre-trained strategy is demonstrated in Section 4.3.
+Methods REVERIE R2R
+Val Unseen Test Unseen Val Unseen Test Unseen
+SR SPL RGSPL SR SPL RGSPL $\downarrow SR SPL NE\downarrow$
+RCM (Wang et al. 2019) 9.2 6.9 3.8 7.8 6.6 3.1 43 - 6.0 43 38 6.1
+GridMM (Wang et al. 2023c) 51.3 36.4 24.5 53.1 36.6 23.4 75 64 - 73 62 -
+KERM (Li et al. 2023b) 49.0 34.8 24.1 52.2 37.4 23.1 71.9 60.9 3.2 69.7 59.2 3.6
+AZHP (Zhan et al. 2024) 49.0 36.2 24.1 52.5 36.1 22.5 71 60 3.2 69 59 3.4
+FDA (He et al. 2024) 47.5 35.9 24.3 49.6 36.4 22.0 72 64 3.4 69 62 3.41
+CONSOLE (Lin et al. 2024) 50.0 34.4 23.3 55.1 37.1 22.2 73 63 3.0 72 61 3.3
+ACME (Wu et al. 2025) 49.5 32.4 24.0 51.9 34.7 23.6 72.8 62.3 3.12 70.4 61.2 3.68
+DUET (Chen et al. 2022b) 46.9 33.7 23.0 52.5 36.0 22.0 72 60 3.3 69 59 3.65
+SUSA (Ours) 51.7 38.8 26.5 54.3 41.5 27.3 73.0 64.8 3.0 72.5 63.8 3.20
+Table 1: Comparison with the state of the art on REVERIE and R2R.
+Bold highlights the best performance in each column.
+$\downarrow$ indicates better performance with lower values.
+“–”: unavailable statistics.
+% is omitted for brevity.
+Contrastive and Imitation Learning.
+During fine-tuning, we employ a contrastive learning loss $\mathcal{L}_{cl}$ to align the hybrid environmental representations $E^{hyb}$ with the corresponding instructions $\mathcal{I}$:
 $\textstyle\mathcal{L}_{cl}=-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{i})/t)}{\sum_{j=1}^{B}\exp((E^{\mathrm{hyb}}_{i},\mathcal{I}_{j})/t)}$ (10)
@@
 where $B$ and $t$ represent the batch size and the temperature.
-Then, we train the model based on hybrid environment representations by performing single-step action prediction [5, 6] with a contrastive learning loss $\mathcal{L}_{cl}$.
-The hybrid single action prediction loss in behavior cloning is computed as $\mathcal{L}_{hsap}^{gt}=\sum_{t=1}^{T}-\log p(a_{t}^{gt}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$ and $\mathcal{L}_{hsap}^{\star}=\sum_{t=1}^{T}-\log p(a_{t}^{\star}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$, where $a_{t}^{gt}$ represents the ground truth action, and $a_{t}^{\star}$ denotes the pseudo-target action sampled from current exploration map.
+Then, we train the model based on hybrid environment representations by performing single-step action prediction (Chen et al.
+2021, 2022b) with a contrastive learning loss $\mathcal{L}_{cl}$.
+The hybrid single action prediction loss in imitation learning is computed as $\mathcal{L}_{hsap}^{gt}=\sum_{t=1}^{T}-\log p(a_{t}^{gt}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$ and $\mathcal{L}_{hsap}^{\star}=\sum_{t=1}^{T}-\log p(a_{t}^{\star}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$, where $a_{t}^{gt}$ represents the ground truth action, and $a_{t}^{\star}$ denotes the pseudo-target action sampled from current exploration map.
 The final training loss is then expressed as follows:
@@
 where $\lambda_{1}$ and $\lambda_{2}$ denote loss weights.
-During the inference phase, the agent is tasked with sequentially navigating to the predicted navigable nodes, following the shortest path algorithm [13], without making any abrupt “jumps” between nodes.
+Method Val Unseen Test Unseen
+SPL RGSPL SPL RGSPL
+GridMM (Wang et al. 2023c) 24.8 3.9 21.2 4.1
+KERM (Li et al. 2023b) 23.1 4.4 - -
+SEAT (Wang et al. 2024b) 24.8 3.9 22.5 4.4
+ACME (Wu et al. 2025) 27.8 4.3 22.3 5.5
+DUET (Chen et al. 2022b) 22.5 3.7 21.4 4.1
+SUSA(Ours) 30.8 6.4 25.4 5.9
+Table 2: Comparison with the state of the art on the SOON.
+Methods Val Unseen Test Unseen
+SR SPL SR SPL
+GridMM (Wang et al. 2023c) 49 41 46 39
+AO-Planner (Chen et al. 2025) 47 33 - -
+ETPNav (An et al. 2024) 57 49 55 48
+DUET (Chen et al. 2022b) 51.9 41.2 47.1 40.5
+SUSA (Ours) 52.7 43.6 50.9 43.9
+Table 3: Comparison with the state of the art on R2R-CE.
+We directly transfer the DUET and SUSA as planners of ETPNav into the continuous environment without pretraining.
+Underline indicates improvements over the baseline.
 4 Experiments
@@
 Datasets.
-We evaluate our model on three diverse VLN benchmarks based on the Matterport3D simulator environment [2].
-R2R [2] focuses solely on navigation following detailed instructions.
-REVERIE [41] requires the agent to recognize the correct object from candidate bounding boxes upon reaching the navigation goal.
-SOON [64] further challenges agents to generate candidate bounding boxes using an object detector, as predefined boxes are not provided.
+We mainly evaluate our model on three diverse VLN benchmarks.
+R2R (Anderson et al.
+2018) focuses solely on navigation following detailed instructions.
+REVERIE (Qi et al.
+2020) requires the agent to recognize the correct object from candidate bounding boxes upon reaching the navigation goal.
+SOON (Zhu et al.
+2021) challenges agents to generate candidate bounding boxes using an object detector.
+Given that discrete agents can serve as planners of ETPNav (An et al.
+2024), we also evaluate SUSA on the continuous R2R-CE (Krantz et al.
+2020).
+R2R-CE is constructed from the discrete Matterport3D and executed in the continuous Habitat simulator (Savva et al.
+2019).
 Metrics.
-We evaluate the navigation performance using standard metrics: Trajectory Length (TL), Success Rate (SR), Navigation Error (NE), and Success Rate weighted by Path Length (SPL).
-Additionally, we use Normalized Dynamic Time Warping (nDTW) and Success weighted by Normalized Dynamic Time Warping (sDTW) to evaluate the fidelity of successful trajectories to their corresponding instructions.
-For the REVERIE and SOON tasks, which involve object identification, we also evaluate object grounding metrics: Remote Grounding Success (RGS), and RGS weighted by Path Length (RGSPL).
+We evaluate the navigation performance using standard metrics: Success Rate (SR), Oracle Success Rate (OSR), Navigation Error (NE), and Success Rate weighted by Path Length (SPL).
+For the REVERIE and SOON tasks, which involve object identification, we also evaluate object grounding metrics: Remote Grounding Success (RGS) and RGS weighted by Path Length (RGSPL).
 Given that SPL optimally balances navigation success rate and trajectory length, we employ it as our primary metric.
 Implementation Details.
-We initialized transformer layers with pretrained LXMERT [48] during pre-training.
 Pre-training was performed for 100k iterations with a batch size of 32.
-Fine-tuning was performed for 25k iterations.
-For fine-tuning, we used batch sizes of 4, 8, and 2 for R2R, REVERIE, and SOON, respectively.
+Fine-tuning was performed for 25k iterations for VLN tasks.
+For fine-tuning, we used batch sizes of 4, 8, 2 and 16 for R2R, REVERIE, SOON and R2R-CE, respectively.
 The default hyperparameters are $\lambda_{1}$ = 0.2 and $\lambda_{2}$= 0.8.
 The node and view dimensions are identical: $d_{v}=d_{n}=d=768$.
-We minimized changes to the baseline (DUET) [6] settings and used no additional annotations.
+We minimized changes to the baseline (DUET) (Chen et al.
+2022b) settings and used no additional annotations.
 All experiments were conducted on a single NVIDIA RTX 4090 GPU.
-Detailed experimental setup is provided in the supplementary material.
 4.2 Comparisons with State of the Art
-We conduct a comprehensive comparison on the REVERIE [41], R2R [2], and SOON [64] datasets.
-For fairness, we exclude methods that involve pre-exploration [66, 46] or utilize large-scale data augmentation [7, 55].
-REVERIE.
-Table 1 presents a comparative analysis on the REVERIE.
-On the test unseen split, SUSA achieves an SPL score of 41.54%, significantly outperforming the previous state-of-the-art (SOTA) method by a large margin of 2.78%.
-Furthermore, we establish new SOTA results for RGS and RGSPL on both the validation and test unseen splits, demonstrating that the introduced textual semantics substantially elevate overall navigation, particularly benefiting object localization performance.
-R2R.
-The compared experiment results on R2R, summarized in  Tab.
-2.
-Compared with the SOTA ESCENE [62] model on the validation and test unseen splits, SUSA achieves 64.85% and 63.83% on SPL, respectively.
-These gains manifest not only an improved navigation performance but also enhanced navigation efficiency, since SPL considers both success rate and trajectory length.
-Methods Validation Unseen Test Unseen
-[…]
-RCM [54] 11.46 43 - 6.09 11.97 43 38 6.12
-HOP [42] 12.27 64 57 3.80 12.68 64 59 3.83
-DSRG [50] - 73 62 3.00 - 72 61 3.33
-PanoGen [28] 13.40 74 64 3.03 14.38 71 61 3.31
-GridMM [56] - 75 64 - - 73 62 -
-FDA [18] 13.68 72 64 3.41 14.76 69 62 3.41
-KERM [31] 13.54 71.95 60.91 3.22 14.60 69.73 59.25 3.61
-AZHP [58] 13.68 71 60 3.25 14.47 69 59 3.43
-CONSOLE [34] 13.59 73 63 3.00 14.31 72 61 3.30
-ESCENE [62] 10.80 68 64 3.39 11.89 66 63 3.77
-DUET [6] 13.94 72 60 3.31 14.73 69 59 3.65
-SUSA (Ours) 12.18 73.05 64.85 3.06 13.27 72.56 63.83 3.23
-Table 2: Comparison with state of the art on the R2R dataset.
-Bold and Underline highlight the best and the runner-up performance in each column.
-$\downarrow$ indicates better performance with lower values.* expressing our reimplementation results.
-SOON.
-On the SOON dataset, as shown in Tab.
-3, SUSA achieves state-of-the-art performance across multiple metrics, including SR, SPL, and RGSPL, attaining SPL scores of 30.82% and 25.47% on the validation and test unseen splits, respectively.
-Moreover, SUSA consistently enhances the baseline (DUET)’s navigation performance and generalization on three datasets, underscoring the effectiveness of integrating textual semantics and depth exploration maps.
-Method Validation Unseen Test Unseen
-[…]
-GridMM [56] 38.92 37.46 24.81 3.91 46.20 36.27 21.25 4.15
-SEAT [53] - 36.87 24.87 3.91 - 35.89 22.55 4.47
-KERM [31] 35.83 38.05 23.16 4.40 - - - -
-DUET [6] 36.20 36.28 22.58 3.75 41.83 33.44 21.42 4.17
-SUSA(Ours) 36.86 42.89 30.82 6.48 36.13 36.87 25.47 5.93
-Table 3: Comparison with the state of the art on the SOON.
+We primarily conduct a comprehensive comparison on three discrete VLN tasks—REVERIE (Qi et al.
+2020), R2R (Anderson et al.
+2018), and SOON (Zhu et al.
+2021).
+For fairness, we exclude methods that involve pre-exploration (Sigurdsson et al.
+2023) or utilize large-scale data augmentation (Wang et al.
+2023b).
+Tables 1 and 2 present a comparative analysis on REVERIE, R2R, and SOON.
+Our method achieves, and in some cases surpasses, the performance of previous state-of-the-art approaches on various metrics.
+For instance, SUSA achieves an SPL/RGSPL of 41.5%/27.3% on the REVERIE test unseen split, significantly outperforming the baseline (DUET) by a large margin of 5.5%/5.3%.
+Furthermore, we briefly transfer the agent into continuous environments to coarsely assess its adaptability and generalization.
+Results in Table 3 indicate that SUSA achieves superior performance over the baseline in the continuous R2R-CE task, highlighting the potential of hierarchical representation enrichment in continuous navigation scenarios.
 4.3 Diagnostic Experiment
-[…]
-ViT $w/o$ 51.07 46.98 33.73 32.15 23.03
-CLIP $w/o$ 57.17 51.55 35.75 35.19 24.40
-CLIP ResNet-50(ImageNet) 54.42 50.18 34.63 34.39 23.66
-CLIP ResNet-50(Gibson) 56.72 52.00 39.21 35.05 26.54
-Table 4: Ablation of different visual representations on the REVERIE validation unseen split. Gibson and ImageNet serve as pre-training sources for the depth encoder in the DSP module.
+RGB Depth SR SPL RGS RGSPL
+ViT $w/o$ 46.9 33.7 32.1 23.0
+CLIP $w/o$ 51.5 35.7 35.1 24.4
+CLIP ImageNet 50.1 34.6 34.3 23.6
+CLIP Gibson 52.0 39.2 35.0 26.5
+Table 4: Ablation of different visual representations on the REVERIE validation unseen split. Gibson and ImageNet serve as pre-training sources for the depth encoder (ResNet-50) in the DSP module.
 $\delta$ REVERIE R2R
-[…]
-$w/o$ 52.00 39.21 35.05 26.54 72.80 62.71 69.33 60.21
-0 53.91 38.87 36.84 26.82 72.20 63.78 70.74 61.46
-0.5 55.07 39.50 37.12 26.70 73.05 64.85 71.05 62.17
-1.0 53.51 37.86 36.69 25.92 72.37 62.95 69.17 60.55
-$adaptive$ 52.43 38.22 36.04 26.24 71.95 63.10 69.98 60.42
-Table 5: Ablation on the relative importance of static and dynamic matching in the TSU module, balanced by the balance factor $\delta$.
+SR SPL RGSPL SR SPL nDTW
+$w/o$ 52.0 39.2 26.5 72.8 62.7 69.3
+0 53.9 38.8 26.8 72.2 63.7 70.7
+0.5 55.0 39.5 26.7 73.0 64.8 71.0
+1.0 53.5 37.8 25.9 72.3 62.9 69.1
+adaptiveadaptive 52.4 38.2 26.2 71.9 63.1 69.9
+Table 5: Ablation of static and dynamic matching strategies in the TSU module, balanced by the balance factor $\delta$.
 1) Image Features vs.
 Structural Features.
-We analyze RGB and depth features contributions within the DSP module (Tab.
-4).
-For RGB images, substituting ViT [11] features with those from CLIP [44] yields marked improvements across all metrics, highlighting that CLIP achieves a better understanding of the environment than its ViT-based baseline counterpart.
-As for depth features, we leveraged depth features extracted by ResNet-50 [17] pre-trained on the ImageNet [9] or Gibson [57] datasets to investigated the effect of depth exploration maps.
+We analyze RGB and depth features contributions within the DSP module (Table  4).
+For RGB images, substituting ViT (Dosovitskiy et al.
+2020) features with those from CLIP (Radford et al.
+2021) yields marked improvements across all metrics.
+As for depth features, we leveraged depth features extracted by ResNet-50 (He et al.
+2016) pre-trained on the ImageNet (Deng et al.
+2009) or Gibson (Xia et al.
+2018) datasets to investigated the effect of depth exploration maps.
 Compared to ImageNet, the Gibson simulator provides spatial structural information that serves as valuable navigation priors.
-Therefore, in  Tab.
-4, the ResNet-50 pre-trained on the Gibson extracts structural features that refine spatial perception, yielding notable improvements in the SPL metric without compromising SR or RGS.
+Therefore, the ResNet-50 pre-trained on the Gibson yields notable improvements in the SPL metric without compromising SR or RGS.
 These improvements indicate that better spatial perception enhances navigation efficiency.
@@
 To assess the impact of static matching ($\delta=0$) and dynamic matching ($\delta=1$) strategies in the TSU module, we perform ablation studies based on balance factor $\delta$ on the R2R and REVERIE.
-Tab.
-5 summarizes these results, where w/o indicates the exclusion of the TSU module, corresponding to the last row results in Tab.
-4.
+Table 5 summarizes these results, where w/o indicates the exclusion of the TSU module.
 Furthermore, adaptive denotes that $\delta$ is a learnable parameter which adaptively adjusts the contribution of each matching strategy.
-Compared to the first row in Tab.
-5, overall, incorporating the TSU module (regardless of static or dynamic matching strategy) improves SR and RGS, albeit at the cost of SPL on REVERIE.
-In contrast, it enhances SPL on R2R.
-While the TSU module can achieve superior performance under the adaptive setting, its performance is inferior when $\delta$ = 0.5.
-Specifically, at $\delta$ = 0.5, the SPL metric reaches 39.50% and 64.85% on the two datasets, with the RGS, nDTW, and sDTW metrics also achieving optimal performance.
-Excellent results demonstrate dual merits of proposed TSU module, which boosts object identification accuracy on the REVERIE task while simultaneously improving the fidelity between navigation trajectories and instructions on the R2R.
-Figure 5: a) and b) show the performance gap ($\downarrow$) between seen and unseen environments, while c) and d) present key metrics ($\uparrow$) for different pretraining strategies on the R2R and REVERIE.
+While the TSU module can achieve superior performance under the adaptive setting, it underperforms when $\delta$ = 0.5.
+Specifically, at $\delta$ = 0.5, the SPL reaches 39.50% and 64.85% on the two datasets.
+Excellent results demonstrate dual merits of TSU module, which boosts object identification accuracy on the REVERIE while simultaneously improving the fidelity between navigation trajectories and instructions on the R2R.
+Figure 5: (a) and (b) show the performance gap ($\downarrow$, lower values indicate better performance.) between seen and unseen environments, while (c) and (d) present key metrics ($\uparrow$) for different pretraining strategies on the R2R and REVERIE.
+ID DSP TSU HAP SR SPL RGSPL
+1 46.9||52.5 33.7||36.0 23.0||22.0
+2 52.0||52.5 39.2||39.5 26.5||24.9
+3 53.1||53.8 38.3||40.8 26.7||26.4
+4 55.0||51.9 39.5||37.8 26.7||25.2
+5 51.7||54.3 38.8||41.5 26.5||27.3
+Table 6: Ablation study on different components of SUSA on the REVERIE (validation||test) unseen split.
 3) Improved Generalization via Pretraining.
-As aforementioned in Sec.
-3.5, pretraining the entire SUSA framework on multiple auxiliary tasks may lead to overfitting, which may negatively impact the generalization in unseen environments, particularly considering the modality differences between textual panorama, depth exploration maps, and RGB representations.
+As aforementioned in Section 3.5, pretraining the entire SUSA framework on multiple auxiliary tasks may lead to overfitting, which negatively impact the generalization in unseen environments.
 As shown in Fig.
-5, we observe that the partial pretraining strategy not only consistently narrow the performance gap between seen and unseen environments but also outperform complete/random pretraining strategies in navigation performance.
-For example, as depicted in Fig.
+5, the partial pretraining strategy not only consistently narrows the performance gap between seen and unseen environments but also outperforms complete/random pretraining strategies in navigation performance.
+In Fig.
 5 (a), the partial pretraining model achieves an SR gap of only 21.11% between the seen and unseen splits on REVERIE, reducing the gap by 5.92% compared to the complete pretraining strategy.
-Additionally, as shown in Fig.
+Additionally, as depicted in Fig.
 5 (c), multiple metrics indicate that partial pretraining results in superior performance.
 Fig.
-5(b) and (d) illustrate similar trends are observed on R2R, further emphasizing the generalization of the partial pretraining strategy.
-ID DSP TSU HRA SR SPL RGSPL
-1 46.98||||52.51 33.73||||36.06 23.03||||22.06
-2 52.00||||52.53 39.21||||39.53 26.54||||24.97
-3 53.17||||53.89 38.31||||40.89 26.73||||26.43
-4 55.07||||51.97 39.50||||37.87 26.70||||25.20
-5 51.75||||54.39 38.86||||41.54 26.56||||27.31
-Table 6: Ablation study on different components of SUSA on the REVERIE (validation||||test) unseen split.
+5 (b) and (d) illustrate similar trends are observed on R2R, further emphasizing the generalization of the partial pretraining strategy.
 4) Overall Design.
-To thoroughly evaluate the efficacy of key components, we conduct ablation experiments on the REVERIE validation and test unseen split in the Tab.
-6.
-As shown in Tab.
-6, an intriguing phenomenon is observed: directly incorporating the DSP and TSU modules (#4) results in an improvement on validation unseen (e.g., SR=55.07, SPL=39.50), while performance on test unseen deteriorates.
-Conversely, contrastive learning within the HRA module (#5) produces opposite outcomes, particularly with respect to the SR metric.
+To thoroughly evaluate the efficacy of key components, we conduct ablation experiments on the REVERIE validation and test unseen splits in Table 6.
+An intriguing phenomenon is observed in Table 6: directly incorporating the DSP and TSU modules (#4) results in an improvement on validation unseen (e.g., SR=55.07, SPL=39.50), while performance on test unseen deteriorates.
+Conversely, contrastive learning within the HAP module (#5) produces opposite outcomes, particularly with respect to the SR metric.
 This may be attributed to the learnable token $v$, which, while pooling hybrid environmental representations, also introduces slight noise that causes fluctuations of navigation performance.
-Nevertheless, the overall performance is progressively improved with the inclusion of DSP, TSU and HRA modules.
-Analysis of the HRA module and the computational efficiency of the entire model is presented in supplementary materials.
+Nevertheless, the overall performance is progressively improved with the inclusion of DSP, TSU and HAP modules.
 4.4 Qualitative Analysis
@@
 6, compared to DUET, our SUSA agent is able to accurately stop near the “living room with a couch.” In contrast, the DUET agent, which relies exclusively on RGB environment representations, is prone to making elusive actions, hindering its ability to accurately follow instructions and reach the target location.
-This manifest that, thanks to the salient landmarks provided by the SUSA architecture, the agent is better able to navigate to the target location by grounding textual instructions.
+This manifests that, thanks to the salient landmarks provided by the SUSA architecture, the agent is better able to navigate to the target location by grounding textual instructions.
 Figure 6: Predicted trajectories of SUSA and DUET on the R2R validation unseen split.
 5 Conclusion
-We present the SUSA architecture, which exploits complementary semantic-spatial representations beyond RGB to facilitate environmental understanding and instruction grounding.
-Specifically, we propose the TSU module, enabling the agent to identify the most instruction-relevant semantic.
-The DSP module enhances spatial awareness by grounding depth exploration maps.
-We further employ contrastive learning to precisely align hybrid environmental representations with instructions.
-Experimental results on the R2R, REVERIE, and SOON benchmarks demonstrate the effectiveness and generalization of our approach.
-Future work will enrich informative environmental representations to advance VLN tasks in continuous environments.
-Supplementary Material
-6 Detailed Experimental Setups
+We present the SUSA architecture, which hierarchical exploits semantic-spatial representations beyond RGB to facilitate environmental understanding and instruction grounding.
+Specifically, we propose the TSU module, enabling the agent to identify the most instruction-relevant textual semantic.
+The DSP module complementarily enhances spatial awareness by incrementally constructing and grounding depth exploration maps.
+Experimental results on the R2R, REVERIE, SOON and R2R-CE benchmarks demonstrate the navigational superiority of our approach.
+Benefiting from hierarchical representation enrichment and alignment, SUSA exhibits excellent generalization performance on unseen environments and modality interpretability.
+6 Acknowledgments
+This work was supported by the National Natural Science Foundation of China (NSFC) under Grants Nos.
+62202139, 62573292 and 62172138, and also supported by the Fundamental Research Funds for the Central Universities under Grants Nos.JZ2025HGTB0226, JZ2024HGTG0310 and Young Elite Scientist Sponsorship Program (No.
+YESS20220409).
+The computations were performed on the High-Performance Computing (HPC) Platform of Hefei University of Technology.
+Appendices
+This appendix provides additional materials to complement the main paper. Specifically:
+  • Section A presents detailed descriptions of metrics, dataset details and pre-training tasks.
+  • Section B offers extended experimental results, including: (1) comparisons with related approaches, (2) analysis of the learned fusion weights across modalities, (3) evaluation of the quality of generated salient landmarks, (4) ablations on the contrastive loss weight, (5) an overview of the overall design on the R2R benchmark, and (6) a discussion on the trade-off between computational efficiency and navigation performance.
+  • Section C showcases additional qualitative examples to analyse the agent’s behavior.
+  • Section D discusses the limitations of our current approach and outlines potential directions for future work.
+Appendix A Detailed Experimental Setups
 Metrics.
 We categorize the metrics into three types: Navigation, Objects Grounding, and Instruction Following.
-Detailed calculations of these metrics can be found in  Tab.
-7.
+Detailed calculations of these metrics can be found in Table 7.
 For Navigation metrics, Trajectory Length (TL): Average path length in meters.
@@
 RGS weighted by Path Length (RGSPL): Prioritizes correct object identification for shorter paths.
-Metric $\uparrow$ $\downarrow$ Definition Dataset
+Metric $\uparrow$ $\downarrow$ Definition
    Navigation:
-TL $\sum_{1\leq i<|P|}d(p_{i},p_{i+1})$ R2R REVERIE SOON
+TL $\sum_{1\leq i<|P|}d(p_{i},p_{i+1})$
 NE $\downarrow$ $d(p_{|P|},g_{|G|})$
@@
    Instruction Following:
-nDTW $\uparrow$ $\exp\left(-\frac{\min_{W\in\mathcal{W}}\sum_{(i_{k},j_{k})\in W}d(p_{i_{k}},q_{j_{k}})}{|P|\cdot d_{th}}\right)$ R2R
-sDTW $\uparrow$ $SR(P,G)*nDTW(P,G)$
+nDTW $\uparrow$ $\exp\left(-\frac{\min_{W\in\mathcal{W}}\sum_{(i_{k},j_{k})\in W}d(p_{i_{k}},q_{j_{k}})}{|P|\cdot d_{th}}\right)$
+sDTW $\uparrow$ SR(P,G)nDTW(P,G)SR(P,G)*nDTW(P,G)
    Objects Grounding:
-RGS $\uparrow$ $\mathbb{I}\Big{(}\max_{j}IoU(B_{i}^{p},B_{j}^{g})\geq 0.5\Big{)}$ REVERIE SOON
+RGS $\uparrow$ $\mathbb{I}\Big(\max_{j}IoU(B_{i}^{p},B_{j}^{g})\geq 0.5\Big)$
 RGSPL $\uparrow$ $\mathrm{RGS}(P,G)\cdot\frac{d(p_{1},g_{|G|})}{\max\{\mathrm{TL}(P),d(p_{1},g_{|G|})\}}$
-Table 7: Common VLN metrics, categorized into three types: Navigation, Instruction Following, and Object Grounding.
-The symbols $\uparrow$ and $\downarrow$ signify if a higher or lower metric value is preferable, respectively.
-P represents the paths taken by the agent, while G stands for the ground truth trajectories.
-The function $\mathbb{I}(\cdot)$ serves as an indicator, with $d(n,P)$ measuring the distance from node $n$ to path $P$ , and $d_{th}$ indicating the distance threshold.
-$W=w_{1..|W|}$ is a warping with $w_{k}=(i_{k},j_{k})\in[1:|P|],[1:|G|]$, respecting the step-size.$B^{p}_{i}$ means the i-th object’s bounding box in agent path (or trajectory).
-Pre-training Tasks.
-Following prior works [6, 5, 30], during pre-training, we employ a suite of multimodal pre-training tasks, including MLM (Masked Language Modeling), MRC (Masked Region Classification), SAP (Single Action Prediction), and OG (Object Grounding).
-OG is specifically designed for object grounding using the REVERIE and SOON datasets.
-MLM and MRC aim to establish a robust vision-language alignment.
-SAP enables the agent to learn preliminary navigation knowledge by predicting the next action during pre-training.
-However, as navigation is inherently a sequential decision-making process, SAP can only learn static, single-step decisions.
-Dataset Train Validation Seen Validation Unseen Test Unseen VLN task
+[…]
+Dataset Train Val Seen Val Unseen Test Unseen VLN task
 instr house instr house instr house instr house
-R2R [2] 14039 61 1021 56 2349 11 4173 18 Fine-grained
-REVERIE [41] 10466 60 1423 46 3521 10 6292 16 Goal-oriented
-SOON [64] 2779 34 113 2 339 5 615 14 Goal-oriented
+R2R (Anderson et al. 2018) 14039 61 1021 56 2349 11 4173 18 Fine-grained
+R2R-CE (Krantz et al. 2020) 10819 61 778 53 1839 11 3408 18 Fine-grained
+REVERIE (Qi et al. 2020) 10466 60 1423 46 3521 10 6292 16 Goal-oriented
+SOON (Zhu et al. 2021) 2779 34 113 2 339 5 615 14 Goal-oriented
 Table 8: Comparison of the three VLN datasets utilized in this study.
 Dataset Details.
-We evaluate the navigation capabilities of the agent on three distinct VLN datasets: the fine-grained R2R [2], as well as the goal-oriented REVERIE [41] and SOON [64] datasets, as shown in Tab.
-8.
+We primarily evaluate the agent’s navigation capabilities on three distinct VLN datasets: the fine-grained R2R (Anderson et al.
+2018), and the goal-oriented REVERIE (Qi et al.
+2020) and SOON (Zhu et al.
+2021).
+To further assess generalization in continuous environments (Savva et al.
+2019), we extend our evaluation to the R2R-CE (Krantz et al.
+2020) task, as shown in Table 8.
 Each dataset is divided into train, validation seen, validation unseen, and test unseen splits.
@@
 The R2R dataset focuses solely on the navigation task, providing detailed path planning for the agent in the instructions.
+In R2R-CE, the average distance between navigable nodes is significantly shorter compared to R2R (0.19 vs.
+2.25 meters), while the number of navigation steps is substantially higher (55 vs.
+4–6).
 SOON is another goal-oriented dataset, containing about 5k instructions, each averaging 47 words in length.
-We evaluated the experimental results of the unseen test split on the public online leaderboards [^1] https://eval.ai/web/challenges/challenge-page/606/participate, https://eval.ai/web/challenges/challenge-page/97/participate.
+Pre-training Tasks.
+Following prior works (Chen et al.
+2022b, 2021), during pre-training, we employ a suite of multimodal pre-training tasks, including MLM (Masked Language Modeling), MRC (Masked Region Classification), SAP (Single Action Prediction), and OG (Object Grounding).
+OG is specifically designed for object grounding using the REVERIE and SOON datasets.
+MLM and MRC aim to establish a robust vision-language alignment.
+SAP enables the agent to learn preliminary navigation knowledge by predicting the next action during pre-training.
+However, as navigation is inherently a sequential decision-making process, SAP can only learn static, single-step decisions.
+When transferring to the R2R-CE task, the agent—serving as part of the planner module—is not trained from scratch in continuous environments, but only fine-tuned.
+We evaluated the experimental results of the unseen test split on the public online leaderboards [^1] https://eval.ai/web/challenges/challenge-page/606/participate, https://eval.ai/web/challenges/challenge-page/97/participate..
 We assure the readers that we will make our code and model checkpoints publicly available.
-7 Additional Experiments
-Validation Unseen
-Methods TL SR SPL RGS RGSPL
-BEVBert [1] - 51.78 36.37 34.71 24.44
-GridMM [56] 23.20 51.37 36.47 34.57 24.56
-SEAT[53] - 49.45 35,51 32.83 23.14
-SUSA(DP) 24.43 52.54 37.59 35.47 25.27
-Depth SUSA(DEM) 22.88 52.00 39.21 35.05 26.54
-KESU [15] 22.93 50.36 35.15 35.02 24.17
-KERM [31] 21.85 50.44 35.38 34.51 24.45
-AACL [33] 22.11 49.42 33.54 33.31 22.49
-CONSOLE [34] - 50.07 34.40 34.05 23.33
-Text SUSA(TP) 23.40 53.17 38.31 37.18 26.73
+Appendix B Additional Experiments
+Val Unseen
+Methods SR SPL RGS RGSPL
+BEVBert 51.78 36.37 34.71 24.44
+GridMM 51.37 36.47 34.57 24.56
+SEAT 49.45 35,51 32.83 23.14
+SUSA(DP) 52.54 37.59 35.47 25.27
+Depth SUSA(DEM) 52.00 39.21 35.05 26.54
+KESU 50.36 35.15 35.02 24.17
+KERM 50.44 35.38 34.51 24.45
+AACL 49.42 33.54 33.31 22.49
+CONSOLE 50.07 34.40 34.05 23.33
+Text SUSA(TP) 53.17 38.31 37.18 26.73
 Table 9: Comparison with related works on the REVERIE validation unseen split.
-7.1 Comparison with Previous Approaches
-We demonstrate the effectiveness of the proposed method through comparisons with related approaches.
-As shown in  Tab.
-9, experiments reveal that directly grounding depth information—whether using depth panoramas (DP) or depth exploration maps (DEM) as input—outperforms prior depth-based methods [1, 56, 53].
+B.1 Comparison with Related Approaches
+We demonstrate the effectiveness of the proposed method through fair comparisons with related approaches.
+As shown in Table 9, experiments reveal that directly grounding depth information—whether using depth panoramas (DP) or depth exploration maps (DEM) as input—outperforms prior depth-based methods (An et al.
+2023; Wang et al.
+2023c, 2024b).
 Notably, using depth exploration maps leads to better navigation performance than depth panoramas.
-Similarly, directly aligning textual semantics with instructions via textual panoramas (TP) surpasses methods [15, 31, 34] that integrate textual semantics into environmental representations.
+Similarly, directly aligning textual semantics with instructions via textual panoramas (TP) surpasses methods (Gao et al.
+2024; Li et al.
+2023b; Lin et al.
+2024) that integrate textual semantics into environmental representations.
 This underscores the superiority of independently grounding each modality, improving the interpretability of each modality’s contribution.
-Figure 7: The weights for hybrid representations in the HRA module change during training (see Eq. 6).
+Figure 7: The weights for hybrid representations in the HAP module change during training (see Eq.  6).
 Figure 8: Similarity matrix for the top 50 most frequent landmarks: REVERIE instructions vs. environments (left), R2R instructions vs. environments (right). Best viewed in color.
-7.2 Fusion Weights of Hybrid Representations.
+B.2 Fusion Weights of Hybrid Representations.
 We quantitatively illustrate the weights (see Eq.
-6) assigned to each branch of the HRA module during training (in Fig.
-7) and inference (in Tab.
-10).
+6) assigned to each branch of the HAP module during training (in Fig.
+7) and inference (in Table 10).
 These weights are normalized between 0 and 1, with higher values denoting stronger dependency on the corresponding environmental representation.
@@
 Despite the increasing dominance of the RGB modality, the agent’s reliance on textual semantics and depth remains above 20%, underscoring their essential role in navigation.
-Tab.
-10 also demonstrates that the agent effectively utilizes the introduced depth spatial and textual semantic information during inference, even though it exhibits a stronger reliance on the RGB modality.
+Table 10 also demonstrates that the agent effectively utilizes the introduced depth spatial and textual semantic information during inference, even though it exhibits a stronger reliance on the RGB modality.
 Splits Local Global
-Semantic $\beta_{1}$ RGB $\beta_{2}$ Depth $\beta_{3}$ RGB $\beta_{4}$
-Validation Seen 0.2147 0.2458 0.2239 0.3156
-Validation Unseen 0.2151 0.2662 0.2216 0.2971
+$\beta_{1}$ $\beta_{2}$ $\beta_{3}$ $\beta_{4}$
+Val Seen 0.2147 0.2458 0.2239 0.3156
+Val Unseen 0.2151 0.2662 0.2216 0.2971
 Table 10: Variation in fusion weights (in Eq.
 6) of hybrid representations during inference on R2R.
-$\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ denotes the fusion weight of textual semantic, local RGB, depth, global RGB representations in the HRA module, respectively.
-7.3 Quality of Generated Salient Landmarks
+$\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ denotes the fusion weight of textual semantic, local RGB, depth, global RGB representations in the HAP module, respectively.
+Figure 9: Analysis of the contrastive loss weight $\lambda_{2}$ in HAP module on R2R validation unseen split.
+B.3 Quality of Generated Salient Landmarks
 The quality of the extracted salient landmarks is crucial for the agent to accurately understand the textual semantics of the environment.
 In Fig.
-8, we compute the similarity between the top-50 most frequently occurring landmarks in our generated salient landmarks and those from R2R [2] and REVERIE [41] to evaluate the quality of the generated salient landmarks.
+8, we compute the similarity between the top-50 most frequently occurring landmarks in our generated salient landmarks and those from R2R (Anderson et al.
+2018) and REVERIE (Qi et al.
+2020) to evaluate the quality of the generated salient landmarks.
 While the types and frequencies of these generated landmarks is different from those in the datasets, they can contain and convey most similar sematic (similarities surpass 0.5).
@@
 Therefore, we use the CLIP text encoder to extract features from the generated salient landmarks.
-Figure 9: Analysis of the contrastive loss weight $\lambda_{2}$ in HRA module on R2R validation unseen split.
-7.4 Analysis of the Contrastive Loss Weight
+[…]
+1 13.41 79.86 71.43 60.63
+2 13.85 80.42 72.80 62.71
+3 13.78 81.65 72.92 63.16
+4 12.18 79.99 73.05 64.85
+5 14.58 82.08 73.95 62.84
+Table 11: Ablation study on different components of SUSA on the R2R validation unseen split.
+B.4 Analysis of the Contrastive Loss Weight
 Our ablation studies indicate that the contrastive learning loss weight, denoted as $\lambda_{2}$ in Eq.
@@
 However, when the contrastive loss weight is too high, performance (SR/OSR) declines, thus overemphasizing contrastive learning loss may lead the agent to prioritize modality alignment over core sequential navigation tasks.
-[…]
-1 13.41 3.30 79.86 71.43 60.63
-2 13.85 3.08 80.42 72.80 62.71
-3 13.78 3.06 81.65 72.92 63.16
-4 12.18 3.06 79.99 73.05 64.85
-5 14.58 3.08 82.08 73.95 62.84
-Table 11: Ablation study on different components of SUSA on the R2R validation unseen split.
-7.5 Overall Design on R2R
+B.5 Overall Design on R2R
 We further supplemented the performance of each key component of SUSA on the R2R task.
-As shown in Tab.
-6, our previous conclusions are reaffirmed: the DSP module allows the agent to better perceive spatial layouts, thus improving navigation efficiency (SPL).
+As shown in Table 11, our previous conclusions are reaffirmed: the DSP module allows the agent to better perceive spatial layouts, thus improving navigation efficiency (SPL).
 The TSU module is better at understanding textual semantics, resulting in higher success rates.
 After integrating the DSP and TSU modules, the agent’s performance improved significantly (e.g., SR=73.05, SPL=64.85).
-When the HRA module is further introduced, the navigation success rate on R2R increases to 73.95, indicating that contrastive learning further enhances instruction-environment alignment.
-However, its SPL metric remains suboptimal, suggesting that the contrastive learning strategy in the HRA module may excel at accurately aligning the environment with instructions but lacks the capability to optimize the navigation path length.
-7.6 Computing Efficiency vs. Navigation Performance
+When the HAP module is further introduced, the navigation success rate on R2R increases to 73.95, indicating that contrastive learning further enhances instruction-environment alignment.
+However, its SPL metric remains suboptimal, suggesting that the contrastive learning strategy in the HAP module may excel at accurately aligning the environment with instructions but lacks the capability to optimize the navigation path length.
+B.6 Computing Efficiency vs. Navigation Performance
 Figure 10: Comparison of computational efficiency and navigation performance on R2R validation unseen split with several popular VLN methods.
-As Fig.
-10 illustrated, we compare the Giga Floating-point Operations (GFLOPs) and navigation performance (i.e., SR/SPL, which are more important metrics) of different VLN models, including DUET [6] , DSRG [50], HAMT [5] and our SUSA.
+As 10 illustrated, we compare the Giga Floating-point Operations (GFLOPs) and navigation performance (i.e, SR/SPL, which are more important metrics) of different VLN models, including DUET (Chen et al.
+2022b) , DSRG (Wang et al.
+2023a), HAMT (Chen et al.
+2021) and our SUSA.
 To ensure a fair comparison across all methods, we performed single-step forward inference with a batch size of 8, an instruction length of 44, and 6 exploration map nodes.
@@
 However, benefiting from the partial pretraining strategy and a shared panorama encoder for both the depth and RGB environmental information, our SUSA architecture notably enhances agent navigation performance with modest computational resources.
-The frozen BLIP-2 and CLIP models provide the agent with off-the-shelf textual semantics and are excluded from the training or inference.
-8 Qualitative Examples
+The frozen BLIP-2 and CLIP models provide the SUSA agent with off-the-shelf textual semantics and are excluded from the training or inference.
+Appendix C Qualitative Examples
 Failure cases.
 We provide an in-depth examination of failure cases, aiming to offer valuable insights for future model improvements.
-Fig.
 11 displays the difference between the navigation trajectory predicted by the proposed SUSA and the ground truth trajectory.
 We found that, although textual semantics can provide the agent with a richer understanding, it may introduces ambiguity in certain scenarios.
-Concretely, when two identical salient landmarks appear in different scenes (e.g., the staircase and bathroom in Fig.
-11), the agent may struggle to distinguish the correct navigable node, leading to erroneous navigation decisions.
+Concretely, when two identical salient landmarks appear in different scenes (e.g., the staircase and bathroom in 11), the agent may struggle to distinguish the correct navigable node, leading to erroneous navigation decisions.
 More Qualitative Examples. We further visualize first-person navigation trajectories on the R2R validation unseen split, as illustrated in Fig. 12.
-Figure 11: Top-down trajectories visualization of failed cases.
-The corresponding instructions are shown at the bottom.
-Partial views of final stopping position and corresponding landmark captions of ground truth and SUSA are also provided
-9 Limitations and Future Work
+Figure 11: Top-down trajectories visualization of failed cases. Partial views of final stop position and corresponding landmark captions are also provided.
+Appendix D Limitations and Future Work
 As previously mentioned, when an agent encounters the same landmark at different navigable points, it may struggle to make confident navigation decisions, leading to failures.
@@
 2) Beyond landmarks, we will construct direction-related textual semantics (e.g., “past,” “turn left”) to better model the spatial relationships between actions in instructions and the environmental layout or historical exploration maps.
-In the future, we also intend to explore the generalizability of our approach, which integrates richer representations from multiple perspectives, in embodied tasks such as continuous navigation and embodied question answering.
+In the future, we also intend to explore the generalizability of our approach, which hierarchically integrates richer representations from multiple perspectives, in embodied tasks such as continuous navigation and embodied question answering.
 Additionally, given the complexity of VLN tasks, current mainstream VLN frameworks are often intricate and cumbersome, with a wide variety of evaluation metrics that are difficult to balance.